### v3

# RoboMemory: A Brain-inspired Multi-memory Agentic Framework for Lifelong Learning in Physical Embodied Systems

Abstract
We present RoboMemory, a brain-inspired multi-memory framework for lifelong learning in physical embodied systems, addressing critical challenges in real-world environments: continuous learning, multi-module memory latency, task correlation capture, and infinite-loop mitigation in closed-loop planning. Grounded in cognitive neuroscience, it integrates four core modules: the Information Preprocessor (thalamus-like), the Lifelong Embodied Memory System (hippocampus-like), the Closed-Loop Planning Module (prefrontal lobe-like), and the Low-Level Executer (cerebellum-like) to enable long-term planning and cumulative learning. The Lifelong Embodied Memory System, central to the framework, alleviates inference speed issues in complex memory frameworks via parallelized updates/retrieval across Spatial, Temporal, Episodic, and Semantic submodules. It incorporates a dynamic Knowledge Graph (KG) and consistent architectural design to enhance memory consistency and scalability. Evaluations on EmbodiedBench show RoboMemory outperforms the open-source baseline (Qwen2.5-VL-72B-Ins) by 25% in average success rate and surpasses the closed-source State-of-the-Art (SOTA) (Claude3.5-Sonnet) by 5%, establishing new SOTA. Ablation studies validate key components (critic, spatial memory, long-term memory), while real-world deployment confirms its lifelong learning capability with significantly improved success rates across repeated tasks. RoboMemory alleviates high latency challenges with scalability, serving as a foundational reference for integrating multi-modal memory systems in physical robots.
Figure 1: The brain-inspired architecture of the RoboMemory resembles the biological nervous system. It maps biological neural components, enabling the agent to interact with diverse environments (real-world, Habitat, ALFRED) and robotic hardware for long-term planning and lifelong learning.
1 Introduction
Driven by the rapid advancements in Vision-Language Models (VLMs) hurst2024gpt ; bai2025qwen2 , VLM-based agents have been increasingly deployed in embodied tasks, leveraging pre-trained knowledge and multimodal understanding capabilities to interact with physical environments park2023generative ; hu2023toward .
Current research primarily focuses on optimizing performance in single tasks within virtual simulators or controlled real-world setups belkhale2024rt ; ahn2022can ; yang2024guiding , or relies on oversimplified memory frameworks. These frameworks lack mechanisms to model interdependencies between distinct tasks. In practice, real-world embodied agents must operate over a lifetime, handling sequential tasks where prior experiences (e.g., learning to “open a fridge” to retrieve an apple) directly shape subsequent performance (e.g., efficiently accessing other objects in similar containers). This demands long-term memory systems that not only retain experiences but also capture cross-task influences, enabling cumulative improvement rather than isolated task execution.
Prior efforts to integrate memory systems into embodied frameworks tan2024cradle ; glocker2025llm ; wang2023voyager ; agashe2024agent ; fu2024msi have advanced long-term planning and lifelong learning. However, most of them concentrate on virtual environments and fail to generalize to the real world. While the embodied frameworks that are designed for the real world yuan2025being ; shi2025hi lack critical memory modules. But on the other hand, a complex memory architecture will suffer from prohibitive latency, which makes it hard to use.
To address these gaps, we propose RoboMemory, a brain-inspired multi-memory framework explicitly designed for lifelong learning in the real world. This highly parallelized, hierarchical architecture enables long-term planning and lifelong learning in real-world environments. Drawing on cognitive neuroscience milner1998cognitive , it comprises four core components (Fig. 1): an Information Preprocessor (thalamus-like) for multimodal integration; a Lifelong Embodied Memory System (hippocampus-like), the core part that adopts a three-tier structure (long-term, short-term, working memory) grounded in cognitive neuroscience to organize experiential data and object spatial relationships, with a parallelizable memory paradigm unifying information updating and retrieval across modules to mitigate latency; a Closed-Loop Planning Module (prefrontal lobe-like) for high-level action sequencing; and a Low-level Executer (cerebellum-like). For robust real-world deployment, RoboMemory employs a dual-system architecture yuan2025being ; shi2025hi where the upper layer (embodied agent) outputs abstract high-level actions, and a Vision-Language-Action (VLA) model with Simultaneous Localization and Mapping (SLAM) system translates these into low-level commands executable by robots.
We validated RoboMemory (excluding the executor) in EmbodiedBench, a long-horizon planning benchmark environment yang2025embodiedbench . Results demonstrate substantial improvements: with Qwen2.5-VL-72b as the backbone, RoboMemory increases average success rates by 25% over the base model and 5% over the closed-source state-of-the-art (SOTA) Claude-3.5-sonnet. Moreover, we test RoboMemory in a real-world environment, where RoboMemory executed 15 diverse tasks twice (one for learning and one for testing). By validating the RoboMemory’s performance between two executions, we verify that the RoboMemory has basic lifelong learning ability in the real world. In addition, we do ablation studies to quantify component-level contributions, and a comprehensive error analysis contextualizes limitations. In conclusion, our contributions fall into three aspects:
  • Inspired by the brain’s unified memory mechanisms, we design a lifelong embodied memory system with four parallel modules (Spatial, Temporal, Episodic, and Semantic) under a unified framework. This framework supports parallelized update and retrieval across modules, mitigating latency accumulation in complex systems while facilitating coherent knowledge integration for lifelong learning.
  • We propose a retrieval-based incremental Knowledge Graph (KG) update algorithm for dynamic spatial memory evolution. It enables efficient, consistent updates by retrieving relevant subgraphs, detecting local conflicts, and merging new information, alleviating scalability bottlenecks of traditional incremental strategies in dynamic environments.
  • RoboMemory enables lifelong learning in real-world physical robots: it executes sequential diverse tasks without memory reset, with experience accumulation driving steady performance improvements. This demonstrates practical long-term autonomous learning in physical scenarios, with reduced reliance on simulated pre-training.
2 Related Work
2.1 VLM/LLM-based Agentic Frameworks in Embodied Tasks
The rapid advancement of VLMs/LLMs has led to diverse agent frameworks in embodied environments yao2022react ; song2023llm ; lin2024swiftsage . Embodied tasks involve partial observability and long-horizon planning, requiring memory systems to retain context. Some use time-ordered context buffers for short-term memory (due to VLMs/LLMs’ limited long-context processing) yao2022react ; packer2023memgpt ; others adopt experience buffers as long-term semantic memory fu2024msi ; shinn2024reflexion . For long-duration tasks, skill libraries serve as procedural memory, with agents accumulating skills via interaction wang2023voyager ; tan2024cradle . Recent efforts integrate diverse memories zhang2023building ; tan2024cradle ; agashe2024agent but focus on virtual/GUI environments, leaving real-world multi-modal memory support for long-term planning under-explored.
2.2 Vision Language Action Model
Current work on VLA models uses imitation learning to output low-level controls from language and visuals black2024pi_0 ; zhao2023learning ; bjorck2025gr00t ; kim2024openvla but is limited to tabletop tasks and single actions, restricting long-horizon planning. VLAs lack long-term execution abilities, while high-level agents excel at planning. Recent works combine high-level frameworks with VLA executors, some augmented with simple memory shi2025hi ; tan2025roboos ; yuan2025being ; yang2025agentic for longer tasks. However, real-world robots need more sophisticated memory to handle continuous multi-task operations over extended periods.
2.3 Memory Frameworks
Many previous works improve long-term planning via memory systems: Voyager wang2023voyager uses a skill library in Minecraft but lacks diverse memory types; CoELA zhang2023building includes procedural, semantic, and episodic memory with a task-specific 2D map; MSI-Agent fu2024msi utilizes insight as long-term memory for in-task learning. Hippo Retrieval Augmented Generation (RAG) gutierrez2024hipporag mimics the hippocampus and introduces KGs as long-term memory indices burgess2002human ; chen2020review , enhancing retrieval. However, the previous approach is mainly focused on constructing a KG with static long context, such as a book, but it is hard to update the graph. We need to update the information in KG for the embodied task. Our approach builds a more general LLM-based memory system using a dynamic KG like Hippo RAG, which is designed for embodied tasks.
Figure 2: RoboMemory architecture with working pipeline and memory mechanisms. (a) Left: The agent’s pipeline. Parallel Step Summarizer and Query Generator in Information Processor (1) generate updates/queries for Lifelong Embodied Memory (2). These memories enable Closed-Loop Planning (3) for tasks like “slice and pick up the apple”—the Planner generates plans, while the Critic and memories adjust decisions via feedback from visual inputs/results (4). (b) Right: Spatial and Semantic memories operate in parallel with isomorphic updates. Internally, Spatial memory maintains a relevance/similarity-updated KG, and Semantic memory manages a Vector DB with analogous logic.
3 RoboMemory
As illustrated in Fig. 2, the key design of RoboMemory is a unified memory paradigm. This paradigm aims to streamline memory operations: during updates, only memory items relevant to new information are targeted, while retrieval leverages rule-based information gathering to enhance efficiency. In the following sections, we detail the specific design of each component and how they integrate within this framework to enable real-world lifelong learning.
3.1 Information Preprocessor
For each time-step $i$, RoboMemory receives a visual observation $\mathcal{O}_{i}$—a single RGB frame in simulation or a short video snippet from the onboard camera on the real robot—capturing what the agent sees while acting. An Information Preprocessor constitutes the system’s perceptual front-end, converting this multimodal input into text that can be indexed and searched. To keep latency low, the preprocessor runs two lightweight modules in parallel:
  1. 1.
    Step summarizer $\mathcal{S}$: Transforms $\mathcal{O}_{i}$ into a concise textual description $s_{i}$ of the just-executed action. The string $s_{i}$ is stored as the system’s working memory.
  2. 2.
    Query generator $\mathcal{Q}$: Derives a query $q_{i}$ from the same observation $\mathcal{O}_{i}$, which is used to probe long-term memory for relevant episodes.
Together, $\mathcal{S}$ and $\mathcal{Q}$ provide a swift, text-based interface between raw sensory data and RoboMemory’s retrieval machinery.
3.2 Lifelong Embodied Memory System
RoboMemory incorporates a lifelong memory system with four modules: Spatial, Temporal, Episodic, and Semantic. The four-module structure is designed to enable continuous learning in dynamic real-world environments. To mitigate latency from repeated VLM invocations, we unify the four modules’ update and retrieval processes into a single paradigm with parallel implementation. As shown in Fig. 2, all four modules perform updates and retrievals in parallel. Thus, the framework avoids latency escalation despite multiple memory components.
Memory modules in Lifelong Embodied Memory System have three types of update frequency:
  1. 1.
    Action-level, which is updated once per action;
  2. 2.
    Task-level, which is updated only after the completion of each task;
  3. 3.
    Mixed-level, which is updated at both the action level and the task level.
Algorithm 1 Retrieval-based Incremental Knowledge Graph Update Algorithm
0: New triplet set $T_{new}$, main knowledge graph $G=(V,E)$, latest retrieval entities $Entity_{l}$, entity embeddings $\mathcal{E}:V\rightarrow\mathbb{R}^{d}$, similarity threshold $\theta$, conflict detection $VLM_{dt}$, retrieval function $\text{Retrieve}(\cdot)$
0: Updated consistent knowledge graph $G^{\prime}$
1: Step 1: Retrieve relevant subgraph based on new triplets
[…]
3: Step 2: Create local working subgraph
[…]
5: Step 3: Conflict detection within local subgraph
[…]
7: Step 4: Generate updated subgraph
[…]
9: Step 5: Merge updated subgraph back to main graph
[…]
11: Update vector database $V$ with new embeddings
12:return  $G^{\prime}$
3.2.1 Spatial-Temporal Memory System
To adapt to dynamic real-world environments, we design the Spatial-Temporal Memory. The spatial and temporal memory update in action-level frequency. The Temporal memory is a first-in-first-out (FIFO) buffer whose size is $N$. The buffer can store $N$ summary steps $s_{[i:i+N]}$. When the buffer is full, it will be fully cleared. Then, we use an LLM to summarize the short-term memories in the discarded $N$ step into a single entity, inserted as the first entry of the buffer.
The Spatial Memory is a dynamically updated KG-based module that addresses LLMs’ limitations in implicitly extracting spatial information from temporal memory. This system dynamically records spatial relationships in the scene while efficiently maintaining a dynamic KG. Unlike previous KG construction algorithms gutierrez2024hipporag ; edge2024local , which are designed for static information gathering. Our spatial memory needs to be updated frequently. To speed up the KG update process, we design a two-phase progressive approach: (1) Rapid Response Phase: New information is quickly buffered to avoid losing observations. (2) Local Integration Phase: On buffer saturation or conflict detection, affected local subgraphs are integrated (entity de-duplication, relationship merging, conflict detection).
This algorithm limits KG updates to the environment. For each update, we only consider relevant segments, controlling efficiency and updater context length for better performance. The algorithm is shown in Algorithm 1. Meanwhile, to illustrate the growth of dynamic spatial memory, we provide a sample in Appendix D.
Our retrieval-based incremental KG update algorithm has provable efficiency guarantees. Specifically, for a KG with $n$ nodes and maximum degree $D$, the number of nodes processed in each update is bounded by $O(D^{K})$ where $K$ is the retrieval hop distance (see Appendix C for formal proof). This ensures scalability even as the spatial memory grows over time.
3.2.2 Lifelong Learning System
In real-world scenarios, agents process sequential tasks over their lifetime, requiring continuous improvement through prior experiences. Drawing on cognitive psychology’s classification of human long-term memory, we divide our system into episodic and semantic memory: the former records agent-environment interaction histories, while the latter extracts experiential insights to support long-term task reasoning. This update process mirrors human daily experience consolidation during sleep maboudi2024retuning .
Episodic Memory: It captures task-level interactions, accounting for temporal interdependencies between sequential tasks in the same environment. The agent must memorize what it has done before to complete future tasks. Moreover, task-level interactions can be a reference that may help the agent improve its plan in the future.
Semantic Memory: It accumulates step-by-step action usage experiences (based on invoked actions and outcomes) to inform action arrangement. Post-task, it summarizes temporal memory, distilling successes from completed tasks and identifying failure causes/improvement strategies from unsuccessful ones, thus enabling both action-level and task-level learning across both task and action levels.
In implementation, both episodic and semantic memory share the same RAG framework consisting of an extractor, updater, and RAG storage (each entry is a memory entity). Post-task, the Extractor summarizes the task’s Spatial-Temporal Memory into a new memory entity. The RAG then retrieves similar existing entities (old information) from the RAG. Then, the Updater deletes, adds, or updates old memory entities according to new information. After that, we write the updated memory entities back to the RAG. Because we only update memory entities similar to new information, efficiency is ensured by restricting updates to old memory entities relevant to the new entity instead of all memory entities stored in the RAG.
3.3 Closed-Loop Planning Module for Dynamic Environment
The Closed-Loop Planning Module integrates information about the current task provided by the Spatial-Temporal Memory, Semantic and Episodic information recorded in long-term memory, and current observations to perform action planning. Each action is planned and passed on to the low-level executor for execution.
To enable closed-loop control in embodied environments, the Closed-Loop Planning Module adopts the Planner-Critic mechanism lei2025clea , which consists of the planner and the critic module. For each planning step, the planner generates a long-term plan consisting of multiple steps. However, due to the dynamics of embodied environments, the action sequence in the long-term plan may become outdated during the execution of the plan. Thus, before executing each step, we use the Critic model to evaluate whether the proposed action in this step remains appropriate under the latest environment. If not, the planner will replan based on the latest information. The demonstration of this process is shown in Fig. 2.
However, our experiments reveal that the original Planner-Critic mechanism may suffer from infinite loops. In the original mechanism, the first step of the action sequence output by the Planner is evaluated by the Critic before execution, which can lead to an infinite loop: if the Critic always demands replanning, no action will ever be executed. To address this, we modified the Planner-Critic mechanism so that the first step is not evaluated by the Critic. This ensures that even if the Critic persistently demands replanning, the RoboMemory will still execute actions.
3.4 Low-level Executer
The RoboMemory framework is a two-layer hierarchical agent framework. This design enables RoboMemory to accomplish longer-term tasks in the real world. The upper layer is responsible only for high-level planning, while the Low-level Executor carries out the actions planned by the upper layer in the real environment.
We use $\pi_{0}$ black2024pi_0 as the action executor in the real world: actions planned by RoboMemory are converted into arm and chassis movements via the Low-level Executer, with LoRA fine-tuning applied to optimize performance in real-world robotic tasks. More details are provided in Appendix B.
Table 1: Comparison of Success Rates (SR) and Goal Condition Success Rates (GC) across difficulty levels between RoboMemory and baseline methods on EB-ALFRED.
Method Type Avg. Base Long
SR GC SR GC SR GC
Single VLM-Agents
GPT-4o Closed-source 59.0 % 68.3 % 64.0 % 74.0 % 54.0 % 62.5 %
GPT-4o-mini 17.0 % 32.4 % 34.0 % 47.8 % 0.0 % 17.0 %
Claude-3.5-Sonnet 62.0 % 63.3 % 72.0 % 72.0 % 52.0 % 54.5 %
Gemini-1.5-Pro 64.0 % 69.7 % 70.0 % 74.3 % 58.0 % 65.0 %
Gemini-2.0-flash 60.0 % 63.9 % 62.0 % 65.7 % 58.0 % 62.0 %
Llama-3.2-90B-Vision-Ins Open-source 27.0 % 33.9 % 38.0 % 43.7 % 16.0 % 24.0 %
InternVL2.5-78B 40.0 % 45.7 % 38.0 % 42.3 % 42.0 % 49.0 %
InternVL3-78B 37.0% - 38.0% - 36.0% -
InternVL2.5-38B 23.3% 31.3% 36.0% 37.3% 26.0% 36.5%
Qwen2.5-VL-72B-Ins 42.0 % - 50.0 % - 34.0 % -
VLM-Agent Frameworks
Voyager (Qwen2.5-VL-72B-Ins) Baselines 44.0% 63.7% 56.0% 73.2% 32.0% 54.2%
Reflexion (Qwen2.5-VL-72B-Ins) 29.0% 43.2% 48.0% 54.0% 10.0% 33.0%
Cradle (Qwen2.5-VL-72B-Ins) 43.0% 54.6% 54.0% 67.9% 32.0% 41.0%
RoboMemory (Qwen2.5-VL-72B-Ins) Ours 67.0 % 78.4 % 68.0 % 75.5 % 66.0 % 81.3 %
4 Experiments
4.1 Experimental Setup
4.1.1 Benchmarks
To evaluate the task planning ability of RoboMemory, we select a subset of the EB-ALFRED benchmark from EmbodiedBenchyang2025embodiedbench . We selected the Base and Long subsets because they aim to test the agent’s planning ability. The Base and Long subset comprises 100 tasks for complex long-term embodied tasks. The EB-ALFRED environment provides a visually-grounded operational setting that closely mimics real-world conditions (see appendix B for environment details), enabling direct comparison with established baselines.
Additionally, we have also evaluated RoboMemory’s capabilities on the EB-Habitat benchmark yang2025embodiedbench . For detailed results, please refer to Appendix A.
Moreover, we set up a real-world environment to test the lifelong learning ability of RoboMemory in the real world.
4.1.2 Settings & Baselines
To facilitate comparisons, we consider two types of baselines. First, we choose the advanced closed-source and open-source VLMs as a single agent. We compare their performance with RoboMemory. For closed source VLMs, we choose GPT-4o and GPT-4o-mini GPT-4o-mini ; hurst2024gpt , Claude3.5-Sonnet Claude-3.5-Sonnet , Gemini-1.5-Pro and Gemini-2.0-flash team2024gemini ; Gemini2.0 . For open source VLMs, we choose LLama-3.2-90B-Vision-Ins llama3.2 , InternVL-2.5-78B/28B chen2024expanding , InternVL-3-72B zhu2025internvl3 , and bai2025qwen2 . Secondly, we choose three agent frameworks: (1) Reflexion shinn2024reflexion , which introduces a simple long-term memory and a self-reflection module. Reflexion uses the self-reflection module to summarize experiences as long-term memory, thereby enhancing the model’s capabilities. (2) Voyager wang2023voyager , which utilizes a skill library as its procedural memory, is a widely used baseline for embodied agent planning. (3) Cradle tan2024cradle , which proposes a general agent framework with episodic and procedural memory and gains good performances at various multi-model agent tasks.
In our experiments, each agent framework is tested using Qwen2.5-VL-72b-Ins qwen2.5 . The Qwen2.5-VL-72b-Ins represents a high-performing open-source alternative. Notably, the Qwen2.5-VL-72b-Ins demonstrates performance comparable to advanced closed-source VLMs in several benchmark tasks white2024livebench . We use the Qwen3-Embedding model zhang2025qwen3 to create embedding vectors for RAGs in RoboMemory. For the Low-level Executor, since EB-ALFRED provides high-level action APIs, we use the low-level executor provided by EmbodiedBench instead of the VLA-based method.
4.1.3 Evaluation metrics
We define two evaluation metrics to assess the performance: (1) Success Rate (SR), which is the ratio of completed tasks to the total number of tasks in each difficulty level. This metric reflects the agent’s ability to complete tasks across randomly generated scenarios. (2) Goal Condition Success Rate (GC), which is the ratio of intermediate conditions achieved to the maximum possible score in each scenario. An GC of 100% indicates that the task is completed in the given scenario. These two metrics can be computed as:
$SR=\mathbb{E}_{x\in\mathcal{X}}\left[\mathds{1}_{SCN_{x}=GCN_{x}}\right]$ (1)
$GC=\mathbb{E}_{x\in\mathcal{X}}\left[\frac{SCN_{x}}{GCN_{x}}\right]$ (2)
Where $\mathcal{X}$ denotes the test subset, and $x$ represents a test task. The success condition number ($SCN_{x}$) refers to the number of conditions the agent has accomplished, while the global condition number ($GCN_{x}$) indicates the total number of conditions required for task completion. The task is considered successful if $SCN_{x}=GCN_{x}$.
Figure 3: The reason why RoboMemory failed to complete the task.
Table 2: Ablation Study on RoboMemory’s Success Rate (SR)
Method Avg. Base Long
RoboMemory 67% 68% 66%
- w/o critic 55 % 60 % 50%
- w/o spatial memory 47 % 52 % 42 %
- w/o long-term memory 57% 66% 48%
4.2 Main results
As shown in Table 1, our model achieves significant improvements over both single VLM agents and Agent frameworks on the EB-ALFRED. Compared to the SOTA Single VLM-Agent model, Claude3.5-Sonnet, RoboMemory with Qwen2.5-VL-72B-Ins backbone improves the average SR by 5% and GC by 15%. This demonstrates RoboMemory’s superiority over single VLM-Agents, proving that an Agent framework with open-source models can outperform closed-source SOTA models. Furthermore, when tested against other VLM-Agent frameworks, RoboMemory also shows substantial gains. This is because, unlike other frameworks, RoboMemory’s brain-like memory system provides embodied models with more accurate and persistent contextual information. Additionally, the Planner-Critic mechanism provides a closed-loop planning ability, which helps the RoboMemory gain better performance in long-term tasks. Because the RoboMemory can detect and try to overcome possible failures. And it is more robust when encountering unexpected situations.
4.3 Ablation Studies
We used the full Base and Long Subset from EB-ALFRED to validate RoboMemory’s effectiveness. We removed each component systematically and observed performance changes across task categories. We use the success rate as our metric. Results are shown in Table 2.
4.3.1 Long-term Memory
Adding long-term memory significantly improved RoboMemory’s success rate. The experiment shows that it enables continuous learning while it attempts to complete tasks. The semantic memory learns low-level skills’ properties, such as in what circumstances an action may fail. The temporal memory records all task attempts (successful/failed), providing valuable experience at the task level, giving insight about how to complete a task successfully. This helps the RoboMemory predict action outcomes and avoid ineffective attempts. This ability indicates that the RoboMemory has lifelong learning capability.
4.3.2 Spatial Memory
Spatial memory is crucial for embodied agents, especially given that current pretrained VLMs have limited spatial understanding ability. Our novel dynamic KG update algorithm enables KG-based spatial memory in dynamic environments. This spatial reasoning helps RoboMemory handle partially observable embodied settings.
4.3.3 Critic Module
Table 2 shows performance without the critic module (55% vs 67% with full system). This drop highlights how the critic’s closed-loop planning adapts to dynamic environments. It helps RoboMemory recover from failures faster and handle unexpected situations better.
4.4 Error Analysis
We summarize the common errors of RoboMemory in the previous experiments. We classify errors into three main types: planning errors, reasoning errors, and perception errors.
The planning errors occur when the planner fails to generate correct actions. The reasoning errors occur when the planner and critic cannot properly process input information (including current observations and memory), even when the input is correct. Perception errors occur when incorrect information is provided to the planner-critic module.
We analyze RoboMemory trajectories for failed tasks. We identify error types based on the above definitions. A single task may contain multiple errors. We calculate the occurrence probability of each error type to show RoboMemory’s strengths and weaknesses. The results are shown in Fig. 3.
We can observe that among all error types, the planning errors are the most common. This means that even though the memory modules can provide comprehensive information about the RoboMemory agent’s previous experience and spatial and temporal memory for the current task, the planner module may still not provide good action plans. This may be due to the capability of the pretrained base model.
The most common perception error is the hallucination error. We can observe that although some hallucinations can be handled by the critic module or memory information, there are still some cases in which the planner ignores all insights from memory and critic and fails to complete the task.
The detailed examples and discussions are provided in Appendix D.
Figure 4: Visualization of the experimental environment.
Figure 5: The improvement of RoboMemory after learning in the real world.
4.5 Real-world Robot Deployment
4.5.1 Environmental Setup
To evaluate RoboMemory’s lifelong learning capability in the real world, we designed a kitchen environment inspired by EB-ALFRED and EB-Habitat. The scene contains 5 navigable points, 8 interactive objects, and over 10 non-interactive (but potentially distracting) items. The environment is shown in Fig. 5. In the real world, we use continuous video recordings during action execution (instead of snapshots after the action execution) as RoboMemory’s input. This provides a more temporally coherent perception. We created three task categories (5 tasks each, totaling 15 tasks) with difficulty matching EB-ALFRED’s Base subset (average oracle trajectory length from 5 to 10). Additional hardware experiment details are in Appendix B.
To test the lifelong learning ability of RoboMemory, we run each task twice without clearing long-term memory between attempts. Success rates for first and second attempts are shown in Fig. 5.
4.5.2 Main results
The second attempt showed significantly higher success rates. This proves RoboMemory’s long-term memory effectively guides subsequent tasks in real embodied environments. Key observations include:
  1. 1.
    Closed-loop error recovery: RoboMemory retries failed actions when possible, even if the low-level executor (VLA model) fails.
  2. 2.
    Spatial reasoning: RoboMemory remembers object locations and spatial relationships using its memory.
  3. 3.
    Lifelong learning: RoboMemory analyzes failure causes reasonably. These analyses guide future decisions.
Detailed examples demonstrating these capabilities and further discussions are provided in Appendix D.
Moreover, we observed a significant drop in task success rates when deploying the agent with the Low-level Executor in real-world environments. This performance degradation primarily stems from the executor’s inherent limitations: (1) The VLA model exhibits unreliable instruction-following capabilities, frequently failing during grasping actions or selecting incorrect objects; (2) Pre-trained VLM models demonstrate inadequate video understanding - while capable of recognizing static objects, they struggle to interpret dynamic visual information such as action failures or state changes. These limitations collectively contribute to the reduced performance compared to simulated environments.
5 Conclusion and Future Work
In conclusion, RoboMemory, a brain-inspired multi-memory framework, enables lifelong learning and long-term planning in real-world embodied systems by addressing key challenges: continuous learning, memory latency, task correlation capture, and planning infinite loops. Experimental results on EmbodiedBench show it outperforms SOTA closed-source VLMs and agent frameworks, with ablation studies validating critical components like the Critic module and spatial/long-term memory. Real-world deployment confirms its lifelong learning capability via improved success rates in repeated tasks. While limited by reasoning errors and executor reliance, it paves the way for generalizable memory-augmented agents, with future work focusing on refining reasoning and enhancing execution robustness.
A key unsolved problem in current hierarchical agent research for embodied tasks, including ours, is about the interaction between high-level agents and low-level executors (e.g., VLAs). Most existing frameworks use language instructions solely as action instructions from high-level agents. However, some action details are hard to describe with language. Other modalities (e.g., vision) can better represent these details (e.g., grasp points). While our work focuses on long-term planning and lifelong learning for agents, future work can consider enhancing generalization within existing frameworks by improving the interaction method between the VLA and the Agent.
Appendix A Additional Experiments
Figure 6: Comparison of Success Rates (SR) and Goal Condition Success Rates (GC) across difficulty levels between RoboBrain and baseline methods on EB-Habitat.
A.1 Evaluate on EB-Habitat
Similar to EB-ALFRED, we also deploy RoboMemory to EB-Habitat. Both EB-Habitat and EB-HALFRED are subsets of EmbodiedBench. We evaluate our model on the Base and Long subsets. Each subset contains 50 different trails. The results are shown in Fig 6.
The results demonstrate that our RoboMemory can adapt well to different environments. It achieves significant improvements over the baseline across various settings. On average, the success rate increases by 24% compared to the SOTA Multi-Agent Method. The goal-conditioned success rate improves by 12%.
These improvements indicate that RoboMemory enhances the agent’s embodied intelligence in various environments. The key factor is its complete memory system.
Table 3: Robot Action Command For different environments
Action Type EB-ALFRED EB-Habitat Real World
Navigate to object/navigation point find(obj) navigate(point) navigate_to(point)
Pick Up Object pick_up(obj) pick(obj) pick_up(obj)
Drop to Ground drop()
Place to Receptacle put_down() place(rec) put_down_to(rec)
Open Object open(obj) open(obj) open(obj)
Close Object close(obj) close(obj) close(obj)
Turn On turn_on(obj) turn_on(obj)
Turn Off turn_off(obj) turn_off(obj)
Slice Object slice(obj)
Task Complete task_complete()
Appendix B Additional Environment settings
B.1 EB-ALFRED and EB-Habitat
We adopt the same environment parameters as in EmbodiedBench. The maximum steps per task are set to 30, with image inputs of size 500 $\times$ 500. The temporal memory buffer length is set to 3. However, we modify the action formats of EB-ALFRED and EB-Habitat to better simulate real-world scenarios. Specifically, we define different action APIs (Python functions), where each action takes an object parameter indicating its target. We extract all possible objects from the environment as inputs to the Agent. The Agent must select appropriate actions and object parameters based on task requirements. Compared to the original interaction method in EmbodiedBench (which enumerates all possible actions, including both action names and target objects, and requires the Agent to choose), our approach offers greater flexibility. The detailed action APIs are presented in Table 3.
Since EB-ALFRED and EB-Habitat provide comprehensive high-level action APIs, we do not employ the VLA-Based Low-Level Executor in these environments. Instead, we utilize the built-in low-level controllers from EmbodiedBench.
B.2 Real-world experiments
We construct a common kitchen scenario to evaluate the RoboMemory framework’s lifelong learning capabilities in real-world settings. Using Mobile ALOHA [14] as our physical robotic platform, we design three categories of tasks: (1) Pick up & put down: The agent must locate a specified object among all possible positions and place it at a designated location. This task tests the model’s basic object-searching and planning abilities. (2) Pick up, operate & put down: Building upon the first task, the agent must additionally perform operations such as heating or cleaning the object. This task requires longer-term planning, which is crucial in embodied environments. (3) Pick up, gather & put down: The agent must place specified objects into a movable container and then move the container to a target location. This task evaluates the agent’s understanding of object relationships, requiring it to remember the positions of at least two objects (the container and the target item) and their spatial relationship.
To adapt to the real-world setup, we define high-level action APIs similar to those in EB-ALFRED and EB-Habitat. Additionally, we train a VLA-based model to execute tasks according to our action APIs. The detailed action APIs are presented in Table 3.
For the low-level executor, we use one main camera and two arm-mounted cameras as input, each with a resolution of 640 $\times$ 480. The temporal memory buffer length is set to 3.
In our experiments, we set the maximum steps per task to 15. We also provide an API for actively terminating tasks. Since real-world environments lack direct success/failure feedback, RoboMemory must autonomously determine task completion. To prevent excessively long task execution, we enforce termination after 15 steps if no success is achieved. A single main camera (640 $\times$ 480 resolution) records video during action execution as input for RoboMemory’s higher-level processing.
Table 4: Dataset statistics and training hyperparameters for robotic manipulation tasks.
Dataset Statistics Training Configuration
Action Type #Episodes Parameter Value
Turning on/off faucet 142 Optimizer AdamW
Picking up & Placing basket on counter 63 Batch size $32\times 6$
Picking up & Placing basket in sink 72 Training steps 10,000
Picking up & Placing banana into basket 114 Learning rate $6.12\times 10^{-5}$
Throwing bottle into trash bin 132 warm up step 500
Placing gum box on dish 120 LoRA Configuration
Picking up & Placing cup on plate 51 rank 16
Picking up & Placing dish into sink 69 $\alpha$ 16
Throwing paper ball into trash bin 135 Resource Usage
Open/close oven 142 GPU A100-80GB $\times$ 6
Total episodes 1040 Training time 12 hours
B.3 Training Details of Low-Level Executer
We use the $\pi_{0}$ model as our foundation model. We collected 1,040 data samples over 10 types of tasks for fine-tuning. We use LoRA [17] fine-tuning to save resources during fine-tuning. The specific fine-tuning parameters and action types are given in Table 4. For tasks involving both pick-up and place actions, we split these tasks into separate pick-up and place actions. These are then treated as two distinct data samples during training. The separation of pick-up and place action allows the VLA to carry an object in its hand. For training, we used a server with six A100-80GB GPUs. The total training time was 12 hours.
Besides, we use the built-in LiDAR SLAM system of the Mobile ALOHA robot base as the navigation action actuator. We define five typical navigation points, similar to EB-Habitat. We used SLAM to navigate between these navigation points.
Appendix C Proof of Dynamic Spatial Memory Update Algorithm
Theorem 1 (Upper Bound on K-hop Node Extraction in Directed Graphs).
Let $G=(V,E)$ be a finite directed graph with maximum out-degree $D\geq 1$, and let $\mathcal{S}\subseteq V$ be a set of $M$ source nodes. Define the K-hop neighborhood $\mathcal{N}_{K}(s)$ of a node $s\in\mathcal{S}$ as the set of nodes reachable from $s$ via directed paths of length at most $K$. Then the total number of distinct nodes in the union of all K-hop neighborhoods,
$\mathcal{N}_{K}(\mathcal{S})=\bigcup_{s\in\mathcal{S}}\mathcal{N}_{K}(s),$
Satisfies the following upper bound:
$|\mathcal{N}_{K}(\mathcal{S})|\leq\begin{cases}M\cdot\dfrac{D^{K+1}-1}{D-1},&\text{if }D>1,\\ M\cdot(K+1),&\text{if }D=1.\end{cases}$
Proof.
For any node $s\in\mathcal{S}$, the number of distinct nodes reachable from $s$ within $i$ hops is at most $D^{i}$, assuming the worst-case scenario where each node encountered has the maximum out-degree $D$, and all neighbors are distinct and non-overlapping.
Thus, the size of the K-hop neighborhood of a single node satisfies:
$|\mathcal{N}_{K}(s)|\leq\sum_{i=0}^{K}D^{i}=\begin{cases}\dfrac{D^{K+1}-1}{D-1},&\text{if }D>1,\\ K+1,&\text{if }D=1.\end{cases}$
Since there are $M$ such source nodes and assuming no overlaps between their K-hop neighborhoods (worst case), the union size satisfies:
$|\mathcal{N}_{K}(\mathcal{S})|\leq M\cdot|\mathcal{N}_{K}(s)|.$
Substituting the bound on $|\mathcal{N}_{K}(s)|$ gives the result. ∎
Theorem 2 (Upper Bound for K-hop Node Extraction in Normalized Directed Graphs).
Let $G=(V,E)$ be a finite directed graph with $|V|=n$ nodes. Assume the maximum out-degree is at most $D_{\max}=Dn$, and the maximum in-degree is at most $N_{\max}=Nn$, where $D,N\in(0,1]$ are constants. Let $\mathcal{S}\subseteq V$ be a set of $M$ source nodes. Define $\mathcal{N}_{K}(\mathcal{S})$ as the union of all nodes reachable from $\mathcal{S}$ via paths of length at most $K$, using only outgoing edges. Then the number of extracted nodes satisfies:
$|\mathcal{N}_{K}(\mathcal{S})|\leq\min\left\{n,\;M\cdot\frac{(Dn)^{K+1}-1}{Dn-1}\right\}.$
In particular, when $Dn\gg 1$, we have the approximation:
$|\mathcal{N}_{K}(\mathcal{S})|\lessapprox M\cdot(Dn)^{K}.$
Proof.
For each node $s\in\mathcal{S}$, the maximum number of reachable nodes within $i$-hops is at most $(Dn)^{i}$ under the assumption of maximum out-degree and no overlap.
Summing over hops from 0 to $K$, we get for each root:
$|\mathcal{N}_{K}(s)|\leq\sum_{i=0}^{K}(Dn)^{i}=\frac{(Dn)^{K+1}-1}{Dn-1}.$
Assuming no overlap among the $M$ source node expansions (worst case), we have:
$|\mathcal{N}_{K}(\mathcal{S})|\leq M\cdot\frac{(Dn)^{K+1}-1}{Dn-1}.$
Since the total number of nodes in the graph is $n$, this quantity is also trivially bounded above by $n$, yielding the result. ∎
Appendix D Supplementary Examples for Qualitative Analysis
Figure 7: Visualization of Spatial Memory’s dynamic update process.
D.1 Example of Dynamic Spatial Memory Update process
In RoboMemory’s Spatial Memory, the KG is dynamically constructed during environment exploration. As illustrated in Fig. 7, we demonstrate the progressive expansion of the KG in Spatial Memory as the agent navigates through the environment. The figure indicates a continuous growth in the number of both nodes and edges of the KG as exploration progresses.
Notably, the KG undergoes dynamic updates through RoboMemory’s environmental interactions. For example, the initial KG state displays the relation “I am near the apple. But as the agent picks up the apple in the third step, in the fourth KG, the relationship becomes “I hold the apple”. This demonstrates RoboMemory’s capability for dynamic KG maintenance and expansion.
By querying this KG, the Planner-Critic module gains access to rich spatial information, empowering RoboMemory with robust spatial memory capabilities that significantly enhance its performance in both TextWorld and EmbodiedBench environments.
D.2 Real World
Figure 8: Case that a task is failed but the experience can help RoboMemory to succeed in the next try.
In Fig. 8, we demonstrate an example of RoboMemory learning through trial and error in a real-world environment. Our task is “place a banana into the oven.” This task required RoboMemory to complete the objectives of finding the banana, picking it up, and transporting it to the oven. We observed that RoboMemory became stuck in an infinite loop during the first attempt. The banana was randomly placed on the “kitchen counter,” but RoboMemory overlooked this navigation target and remained trapped exploring other navigation targets instead.
However, based on this bad attempt, the semantic memory summarized that the robot should not repeatedly search in locations where the “banana” could not be found. Meanwhile, the episodic memory recorded what RoboMemory had done and the outcomes during the first attempt. Based on the information provided by semantic and episodic memory, in the second attempt, RoboMemory recognized that it had not previously tried navigating to the “kitchen counter.” After attempting this, it successfully completed the task. This example illustrates the role of RoboMemory’s long-term memory.
Figure 9: Case that a task is successful.
We also provide an example that completes the task in the first attempt. The example is shown in Fig. 9. This example demonstrates that the RoboMemory has the ability to handle some relatively complex tasks in the real world. The task in this example is “Place a box of gum into the basket and put the basket on the kitchen counter”. Because two objects in different positions are involved in this task, RoboMemory has to memorize the position of at least one object to achieve the goal. With the help of the spatial memory, RoboMemory completes the task successfully.
D.3 EB-ALFRED
We select three examples in EB-ALFRED to show the errors that RoboMemory may encounter and the reasons why or why not RoboMemory can achieve the goal.
D.3.1 Successful example
Figure 10: Case that a task is successful with the help of the critic and spatial memory modules.
We select a successful example to show how RoboMemory performed in the EB-ALFRED environment. The example trajectory is shown in Fig. 10.
The task of this example is “set a plate with a spoon on it on the kitchen table”. However, in step 10, the Planner seems to ignore the temporal information from memory modules. RoboMemory thinks that it still needs to pick up the spoon (even though it has already placed a spoon in the plate). However, with the help of the critic, it finally becomes aware that picking up another spoon is redundant, so RoboMemory goes back to the current trajectory and successfully completes the task at the end.
In this example, RoboMemory successfully overcame the hallucination and eventually achieved the goal. This example demonstrates that the critic module can help RoboMemory to overcome error cases.
D.3.2 Failed example
Figure 11: Case that a task fails in an infinite loop because the critic module failed to stop the agent when its planned action is no longer suitable.
We demonstrate a representative example of the Critical Error. The example trajectory is shown in Fig. 11. In this example, the task involves slicing and heating a tomato and moving the heated tomato slice to the trash can. Initially, RoboMemory successfully sliced the tomato with a knife. But when the planner plans the whole sequence, it forgets to drop the knife before picking up the tomato (this is necessary because in EB-ALFRED, the robot can only hold one object at a time). The critic and the planner should notice this situation and ask the critic to replan, as RoboMemory failed to pick up a tomato slice. However, the critic module ignores this issue, and thus, after it heats the knife instead of a tomato slice, it stacks in an infinite loop.
Figure 12: Case that a task fails in an infinite loop because of inaccurate action planning.
Besides, we provide another example demonstrating a representative failure caused by inaccurate action planning. The example trajectory is shown in Fig. 12. In the trajectory, RoboMemory is asked to place two CDs into the drawer. However, at step 6, the robot failed to select correct CD object. In this experiment, RoboMemory has already put CD_2 into the drawer, but it keeps picking up CD_2 even though the memory has clearly indicated that CD_2 has already been put down. So we classify this as inaccurate action error. This indicates that the planner failed to comprehensively integrate information from both the memory and information-gathering modules, resulting in inaccurate action planning.

### v4

@@
-RoboMemory: A Brain-inspired Multi-memory Agentic Framework for Lifelong Learning in Physical Embodied Systems
+RoboMemory: A Brain-inspired Multi-memory Agentic Framework for Interactive Environmental Learning in Physical Embodied Systems
 Abstract
-We present RoboMemory, a brain-inspired multi-memory framework for lifelong learning in physical embodied systems, addressing critical challenges in real-world environments: continuous learning, multi-module memory latency, task correlation capture, and infinite-loop mitigation in closed-loop planning.
-Grounded in cognitive neuroscience, it integrates four core modules: the Information Preprocessor (thalamus-like), the Lifelong Embodied Memory System (hippocampus-like), the Closed-Loop Planning Module (prefrontal lobe-like), and the Low-Level Executer (cerebellum-like) to enable long-term planning and cumulative learning.
-The Lifelong Embodied Memory System, central to the framework, alleviates inference speed issues in complex memory frameworks via parallelized updates/retrieval across Spatial, Temporal, Episodic, and Semantic submodules.
-It incorporates a dynamic Knowledge Graph (KG) and consistent architectural design to enhance memory consistency and scalability.
-Evaluations on EmbodiedBench show RoboMemory outperforms the open-source baseline (Qwen2.5-VL-72B-Ins) by 25% in average success rate and surpasses the closed-source State-of-the-Art (SOTA) (Claude3.5-Sonnet) by 5%, establishing new SOTA.
-Ablation studies validate key components (critic, spatial memory, long-term memory), while real-world deployment confirms its lifelong learning capability with significantly improved success rates across repeated tasks.
-RoboMemory alleviates high latency challenges with scalability, serving as a foundational reference for integrating multi-modal memory systems in physical robots.
-Figure 1: The brain-inspired architecture of the RoboMemory resembles the biological nervous system.
-It maps biological neural components, enabling the agent to interact with diverse environments (real-world, Habitat, ALFRED) and robotic hardware for long-term planning and lifelong learning.
+Embodied agents face persistent challenges in real-world environments, including partial observability, limited spatial reasoning, and high-latency multi-memory integration.
+We present RoboMemory, a brain-inspired framework that unifies Spatial, Temporal, Episodic, and Semantic memory under a parallelized architecture for efficient long-horizon planning and interactive environmental learning.
+A dynamic spatial knowledge graph (KG) ensures scalable and consistent memory updates, while a closed-loop planner with a critic module supports adaptive decision-making in dynamic settings.
+Experiments on EmbodiedBench show that RoboMemory, built on Qwen2.5-VL-72B-Ins, improves average success rates by 25% over its baseline and exceeds the closed-source state-of-the-art (SOTA) Gemini-1.5-Pro by 3%.
+Real-world trials further confirm its capacity for cumulative learning, with performance improving across repeated tasks.
+These results highlight RoboMemory as a scalable foundation for memory-augmented embodied intelligence, bridging the gap between cognitive neuroscience and robotic autonomy.
+Figure 1: RoboMemory adopts a brain-inspired architecture that maps neural components to agent modules, enabling long-term planning and interactive learning across diverse environments (real-world, Habitat, ALFRED) and robotic hardware.
 1 Introduction
-Driven by the rapid advancements in Vision-Language Models (VLMs) hurst2024gpt ; bai2025qwen2 , VLM-based agents have been increasingly deployed in embodied tasks, leveraging pre-trained knowledge and multimodal understanding capabilities to interact with physical environments park2023generative ; hu2023toward .
-Current research primarily focuses on optimizing performance in single tasks within virtual simulators or controlled real-world setups belkhale2024rt ; ahn2022can ; yang2024guiding , or relies on oversimplified memory frameworks.
-These frameworks lack mechanisms to model interdependencies between distinct tasks.
-In practice, real-world embodied agents must operate over a lifetime, handling sequential tasks where prior experiences (e.g., learning to “open a fridge” to retrieve an apple) directly shape subsequent performance (e.g., efficiently accessing other objects in similar containers).
-This demands long-term memory systems that not only retain experiences but also capture cross-task influences, enabling cumulative improvement rather than isolated task execution.
-Prior efforts to integrate memory systems into embodied frameworks tan2024cradle ; glocker2025llm ; wang2023voyager ; agashe2024agent ; fu2024msi have advanced long-term planning and lifelong learning.
-However, most of them concentrate on virtual environments and fail to generalize to the real world.
-While the embodied frameworks that are designed for the real world yuan2025being ; shi2025hi lack critical memory modules.
-But on the other hand, a complex memory architecture will suffer from prohibitive latency, which makes it hard to use.
-To address these gaps, we propose RoboMemory, a brain-inspired multi-memory framework explicitly designed for lifelong learning in the real world.
-This highly parallelized, hierarchical architecture enables long-term planning and lifelong learning in real-world environments.
-Drawing on cognitive neuroscience milner1998cognitive , it comprises four core components (Fig.
-1): an Information Preprocessor (thalamus-like) for multimodal integration; a Lifelong Embodied Memory System (hippocampus-like), the core part that adopts a three-tier structure (long-term, short-term, working memory) grounded in cognitive neuroscience to organize experiential data and object spatial relationships, with a parallelizable memory paradigm unifying information updating and retrieval across modules to mitigate latency; a Closed-Loop Planning Module (prefrontal lobe-like) for high-level action sequencing; and a Low-level Executer (cerebellum-like).
-For robust real-world deployment, RoboMemory employs a dual-system architecture yuan2025being ; shi2025hi where the upper layer (embodied agent) outputs abstract high-level actions, and a Vision-Language-Action (VLA) model with Simultaneous Localization and Mapping (SLAM) system translates these into low-level commands executable by robots.
-We validated RoboMemory (excluding the executor) in EmbodiedBench, a long-horizon planning benchmark environment yang2025embodiedbench .
-Results demonstrate substantial improvements: with Qwen2.5-VL-72b as the backbone, RoboMemory increases average success rates by 25% over the base model and 5% over the closed-source state-of-the-art (SOTA) Claude-3.5-sonnet.
-Moreover, we test RoboMemory in a real-world environment, where RoboMemory executed 15 diverse tasks twice (one for learning and one for testing).
-By validating the RoboMemory’s performance between two executions, we verify that the RoboMemory has basic lifelong learning ability in the real world.
-In addition, we do ablation studies to quantify component-level contributions, and a comprehensive error analysis contextualizes limitations.
-In conclusion, our contributions fall into three aspects:
-  • Inspired by the brain’s unified memory mechanisms, we design a lifelong embodied memory system with four parallel modules (Spatial, Temporal, Episodic, and Semantic) under a unified framework.
-This framework supports parallelized update and retrieval across modules, mitigating latency accumulation in complex systems while facilitating coherent knowledge integration for lifelong learning.
-  • We propose a retrieval-based incremental Knowledge Graph (KG) update algorithm for dynamic spatial memory evolution.
-It enables efficient, consistent updates by retrieving relevant subgraphs, detecting local conflicts, and merging new information, alleviating scalability bottlenecks of traditional incremental strategies in dynamic environments.
-  • RoboMemory enables lifelong learning in real-world physical robots: it executes sequential diverse tasks without memory reset, with experience accumulation driving steady performance improvements.
-This demonstrates practical long-term autonomous learning in physical scenarios, with reduced reliance on simulated pre-training.
-2 Related Work
-2.1 VLM/LLM-based Agentic Frameworks in Embodied Tasks
-The rapid advancement of VLMs/LLMs has led to diverse agent frameworks in embodied environments yao2022react ; song2023llm ; lin2024swiftsage .
-Embodied tasks involve partial observability and long-horizon planning, requiring memory systems to retain context.
-Some use time-ordered context buffers for short-term memory (due to VLMs/LLMs’ limited long-context processing) yao2022react ; packer2023memgpt ; others adopt experience buffers as long-term semantic memory fu2024msi ; shinn2024reflexion .
-For long-duration tasks, skill libraries serve as procedural memory, with agents accumulating skills via interaction wang2023voyager ; tan2024cradle .
-Recent efforts integrate diverse memories zhang2023building ; tan2024cradle ; agashe2024agent but focus on virtual/GUI environments, leaving real-world multi-modal memory support for long-term planning under-explored.
-2.2 Vision Language Action Model
-Current work on VLA models uses imitation learning to output low-level controls from language and visuals black2024pi_0 ; zhao2023learning ; bjorck2025gr00t ; kim2024openvla but is limited to tabletop tasks and single actions, restricting long-horizon planning.
-VLAs lack long-term execution abilities, while high-level agents excel at planning.
-Recent works combine high-level frameworks with VLA executors, some augmented with simple memory shi2025hi ; tan2025roboos ; yuan2025being ; yang2025agentic for longer tasks.
-However, real-world robots need more sophisticated memory to handle continuous multi-task operations over extended periods.
-2.3 Memory Frameworks
-Many previous works improve long-term planning via memory systems: Voyager wang2023voyager uses a skill library in Minecraft but lacks diverse memory types; CoELA zhang2023building includes procedural, semantic, and episodic memory with a task-specific 2D map; MSI-Agent fu2024msi utilizes insight as long-term memory for in-task learning.
-Hippo Retrieval Augmented Generation (RAG) gutierrez2024hipporag mimics the hippocampus and introduces KGs as long-term memory indices burgess2002human ; chen2020review , enhancing retrieval.
-However, the previous approach is mainly focused on constructing a KG with static long context, such as a book, but it is hard to update the graph.
-We need to update the information in KG for the embodied task.
-Our approach builds a more general LLM-based memory system using a dynamic KG like Hippo RAG, which is designed for embodied tasks.
-Figure 2: RoboMemory architecture with working pipeline and memory mechanisms.
-(a) Left: The agent’s pipeline.
-Parallel Step Summarizer and Query Generator in Information Processor (1) generate updates/queries for Lifelong Embodied Memory (2).
-These memories enable Closed-Loop Planning (3) for tasks like “slice and pick up the apple”—the Planner generates plans, while the Critic and memories adjust decisions via feedback from visual inputs/results (4).
-(b) Right: Spatial and Semantic memories operate in parallel with isomorphic updates.
-Internally, Spatial memory maintains a relevance/similarity-updated KG, and Semantic memory manages a Vector DB with analogous logic.
+Recent advances in Vision-Language Models (VLMs) (hurst2024gpt, ; bai2025qwen2, ) have enabled their growing use in embodied tasks (park2023generative, ; hu2023toward, ).
+VLM-based embodied agents can process multimodal inputs and generate high-level textual commands (e.g., “Pick up the cup”), which require translation via tool APIs to become executable robot actions.
+In contrast, Vision-Language-Action models (VLAs) (kim2024openvla, ; black2024pi_0, ; bjorck2025gr00t, ; chi2023diffusion, ) produce low-level control signals directly but generally rely only on the latest observation.
+This limits their ability to perform long-horizon, multi-step tasks that require reasoning over task history.
+In summary, VLA models enable direct robot control but lack high-level planning capabilities, and VLM-based embodied agents support strategic planning but struggle with direct motor control.
+This highlights a key gap inherent in two distinct technical approaches to embodied intelligence.
+To bridge this gap, recent work (yuan2025being, ; shi2025hi, ; tan2025roboos, ) proposes a “VLM planner + VLA executor” paradigm.
+Here, VLM-based embodied agents serve as high-level planners that decompose complex tasks (e.g., “make a coffee”) into executable sub-instructions (e.g., “grasp the cup”) that VLAs can complete.
+Although this paradigm improves performance on multi-step tasks, prior work suffers from two key limitations in real-world settings.
+First, real-world tasks (e.g., kitchen operations) require navigating across multiple locations to gather objects and tools, but the environment remains only partially observable at any time due to robots’ limited field of view and dynamic occlusions.
+This necessitates a planner with robust spatial awareness and long-term memory to maintain a consistent spatial awareness across viewpoints.
+However, most VLM-based agents rely on chat-style context windows (e.g., logging instruction – feedback pairs (yao2022react, )), which lack mechanisms for maintaining an overview of the environment’s spatial layout.
+Consequently, agents cannot reliably track object locations or recognize previously visited states.
+Second, pretrained VLMs are rarely trained on embodied planning trajectories, especially long-horizon, spatially grounded ones.
+So VLM-based agents often struggle to generalize to real-world settings (yang2025embodiedbench, ).
+To overcome these challenges, VLM-based planners must support interactive environmental learning — the ability to acquire, integrate, and retrieve spatial, episodic, and semantic knowledge during task execution, thereby enabling adaptation through experience.
+To support long-horizon planning and interactive learning in real-world settings, agents require a comprehensive memory system with multiple specialized modules.
+Recent frameworks (tan2024cradle, ; glocker2025llm, ; wang2023voyager, ; agashe2024agent, ; fu2024msi, ; zhao2024expel, ; chen2024automanual, ) have integrated Retrieval-Augmented Generation (RAG)-based memory to enhance planning and interactive environmental learning (gao2023retrieval, ), but most are designed for simulated environments.
+A key limitation for real-world deployment is the absence of spatial memory, which is critical for building spatial awareness and providing context for planning.
+Additionally, existing multi-module memory systems often incur significant inference latency.
+To overcome these limitations, especially the need for memory that is efficient, spatially grounded, and persistent in dynamic environments, we return to the essence of intelligence — how does the human brain plan, remember, and learn in dynamic environments?
+Inspired by cognitive neuroscience, we have designed RoboMemory, a parallel multi-memory architecture that simulates key functional regions of the brain.
+RoboMemory features a hierarchical and parallelized architecture enabling long-term planning and continuous adaptation.
+Drawing inspiration from cognitive neuroscience (milner1998cognitive, ), RoboMemory comprises four core components (Figure 1): (1) Information Preprocessor (thalamus-inspired) for multimodal sensory integration.
+(2) Comprehensive Embodied Memory System (hippocampus-inspired), which organizes experiential and spatial knowledge through a three-tier structure (long-term, short-term, and sensory memory).
+Within this tiered system, four memory modules: Spatial, Temporal, Episodic, and Semantic operate under a unified, parallel-update paradigm to enable coherent knowledge integration while minimizing latency.
+(3) Closed-Loop Planning Module (prefrontal cortex-inspired) for high-level action sequencing.
+These three modules provide a high-level planner with comprehensive sensory and memorization ability.
+(4) Low-level Executor (cerebellum-inspired), consisting of a VLA-based operation model and a SLAM-based navigation model.
+The Low-level Executor directly controls the robot with low-level control signals to navigate and operate in the real-world environment.
+To verify whether RoboMemory truly addresses the problems of long-horizon planning and interactive learning, we evaluate RoboMemory on EmbodiedBench, a long-horizon planning benchmark (yang2025embodiedbench, ).
+Using Qwen2.5-VL-72B as the base model, RoboMemory improves average success rates by 25% over its base model and 3% over the closed-source state-of-the-art model, Gemini-1.5-Pro.
+In real-world trials, RoboMemory executed diverse tasks twice consecutively: once for environmental familiarization (learning phase) and once for memory-augmented execution (testing phase) without resetting memory.
+The observed performance improvement validates RoboMemory’s capacity for interactive environmental learning.
+We further conduct ablation studies and error analysis to quantify component contributions and identify remaining limitations.
+We summarize our contribution as follows:
+  • We propose a brain-inspired unified embodied memory system, integrating four concurrently updated modules (Spatial, Temporal, Episodic, Semantic) into a single framework.
+It enables efficient, comprehensive memory operations and coherent knowledge integration, which are critical for interactive environmental learning in real-world embodied scenarios.
+  • We design a retrieval-based incremental update algorithm for real-time evolution of Spatial Knowledge Graphs (KGs).
+By retrieving relevant subgraphs, detecting local inconsistencies, and merging new observations, it ensures efficient, consistent KG maintenance and addresses the scalability bottleneck of previous KG-based methods in embodied settings.
+  • RoboMemory supports interactive environmental learning for real-world physical robots: it enables sequential diverse tasks without memory reset, with experience accumulation driving steady performance improvements, demonstrating practical long-term autonomous learning in physical scenarios.
 2 RoboMemory
-As illustrated in Fig.
-2, the key design of RoboMemory is a unified memory paradigm.
-This paradigm aims to streamline memory operations: during updates, only memory items relevant to new information are targeted, while retrieval leverages rule-based information gathering to enhance efficiency.
-In the following sections, we detail the specific design of each component and how they integrate within this framework to enable real-world lifelong learning.
+RoboMemory is a hierarchical embodied agent system that equips robots with three core memory capabilities: historical interaction logs, dynamically updated spatial layouts, and accumulated task knowledge.
+As illustrated in Figure 2, each iteration, RoboMemory follows a process of “Perception – Memory – Retrieval – Planning – Execution” process, ensuring that the agent continuously calibrates its memory and behavior in dynamic environments.
+First, the information preprocessor converts multimodal sensor inputs into a textual summary of the current scene, which serves as the primary input to the Comprehensive Embodied Memory.
+Next, the Comprehensive Embodied Memory updates its internal representations, including action histories, object locations, and experiential knowledge.
+After information is updated, the memory system retrieves contextually relevant entries to inform the Closed-Loop Planning Module.
+Then, leveraging this contextual memory, the Closed-Loop Planning Module generates high-level, text-based action instructions.
+Finally, these commands are dispatched to low-level executors, who will directly control the robot and complete the instructions.
+Figure 2: RoboMemory architecture.
+(a) Left: Parallel Step Summarizer and Query Generator generate updates/queries for Comprehensive Embodied Memory.
+These memories enable Closed-Loop Planning for tasks like “slice and pick up the apple”—the Planner generates plans, while the Critic and memories adjust decisions via feedback from visual inputs/results.
+(b) Right: Spatial memory maintains a relevance/similarity-updated KG, and Semantic/Episodic memory manages a Vector DB with analogous logic.
+Besides, Temporal memory is implemented as a linear FIFO buffer that stores step-wise summaries generated by the Step Summarizer.
 2.1 Information Preprocessor
-For each time-step $i$, RoboMemory receives a visual observation $\mathcal{O}_{i}$—a single RGB frame in simulation or a short video snippet from the onboard camera on the real robot—capturing what the agent sees while acting.
-An Information Preprocessor constitutes the system’s perceptual front-end, converting this multimodal input into text that can be indexed and searched.
-To keep latency low, the preprocessor runs two lightweight modules in parallel:
-  1. 1.
-    Step summarizer $\mathcal{S}$: Transforms $\mathcal{O}_{i}$ into a concise textual description $s_{i}$ of the just-executed action. The string $s_{i}$ is stored as the system’s working memory.
-  2. 2.
-    Query generator $\mathcal{Q}$: Derives a query $q_{i}$ from the same observation $\mathcal{O}_{i}$, which is used to probe long-term memory for relevant episodes.
-Together, $\mathcal{S}$ and $\mathcal{Q}$ provide a swift, text-based interface between raw sensory data and RoboMemory’s retrieval machinery.
-3.2 Lifelong Embodied Memory System
-RoboMemory incorporates a lifelong memory system with four modules: Spatial, Temporal, Episodic, and Semantic.
-The four-module structure is designed to enable continuous learning in dynamic real-world environments.
-To mitigate latency from repeated VLM invocations, we unify the four modules’ update and retrieval processes into a single paradigm with parallel implementation.
-As shown in Fig.
-2, all four modules perform updates and retrievals in parallel.
-Thus, the framework avoids latency escalation despite multiple memory components.
-Memory modules in Lifelong Embodied Memory System have three types of update frequency:
-  1. 1.
-    Action-level, which is updated once per action;
-  2. 2.
-    Task-level, which is updated only after the completion of each task;
-  3. 3.
-    Mixed-level, which is updated at both the action level and the task level.
-Algorithm 1 Retrieval-based Incremental Knowledge Graph Update Algorithm
-0: New triplet set $T_{new}$, main knowledge graph $G=(V,E)$, latest retrieval entities $Entity_{l}$, entity embeddings $\mathcal{E}:V\rightarrow\mathbb{R}^{d}$, similarity threshold $\theta$, conflict detection $VLM_{dt}$, retrieval function $\text{Retrieve}(\cdot)$
-0: Updated consistent knowledge graph $G^{\prime}$
-1: Step 1: Retrieve relevant subgraph based on new triplets
-[…]
-3: Step 2: Create local working subgraph
-[…]
-5: Step 3: Conflict detection within local subgraph
-[…]
-7: Step 4: Generate updated subgraph
-[…]
-9: Step 5: Merge updated subgraph back to main graph
-[…]
-11: Update vector database $V$ with new embeddings
-12:return  $G^{\prime}$
+At each time step $i$, RoboMemory receives a visual observation $\mathcal{O}_{i}$: an RGB frame (in simulation) or a short video clip (on physical robots), representing the agent’s observations.
+Since raw visual data is unsuitable for direct use in memory construct and retrieval, RoboMemory first employs an information preprocessor to convert multimodal observations into textual representations, thereby providing a semantic interface for subsequent memory and planning modules.
+The information preprocessor executes two Vision-Language Models (VLMs) in parallel: (1) Step summarizer $\mathcal{S}$: Transforms $\mathcal{O}_{i}$ into a concise textual description $s_{i}$ of the just-executed action.
+The string $s_{i}$ is stored in the system’s working memory.
+(2) Query generator $\mathcal{Q}$: Derives a query $q_{i}$ from the same observation $\mathcal{O}_{i}$, which is used to probe long-term memory for relevant episodes.
+Together, $\mathcal{S}$ and $\mathcal{Q}$ provide a swift, text-based interface between raw sensory data and provide basic information in each iteration for RoboMemory’s Comprehensive Embodied Memory System.
+2.2 Comprehensive Embodied Memory System
+We present a high-level memory system designed to enhance embodied agents’ long-horizon planning and enable interactive learning in dynamic, partially observable environments.
+Built on RAG and KG technologies (gutierrez2024hipporag, ), the system comprises Spatial-Temporal Memory and Long-Term Memory.
+Spatial-Temporal Memory continuously encodes object relationships and environmental layouts, as well as detailed task execution traces.
+These records inform Long-Term Memory, which goes through spatial-temporal information collected in each task, capturing what strategies succeeded or failed and relevant environmental facts (Figure 2).
+The Comprehensive Embodied Memory module receives summarized sensory information from the information preprocessor to update both memory components in real time.
+Critically, all updates and retrievals are performed in parallel, ensuring efficient operation even under multiple active memory modules.
 2.2.1 Spatial-Temporal Memory System
-To adapt to dynamic real-world environments, we design the Spatial-Temporal Memory.
-The spatial and temporal memory update in action-level frequency.
-The Temporal memory is a first-in-first-out (FIFO) buffer whose size is $N$.
-The buffer can store $N$ summary steps $s_{[i:i+N]}$.
-When the buffer is full, it will be fully cleared.
-Then, we use an LLM to summarize the short-term memories in the discarded $N$ step into a single entity, inserted as the first entry of the buffer.
-The Spatial Memory is a dynamically updated KG-based module that addresses LLMs’ limitations in implicitly extracting spatial information from temporal memory.
-This system dynamically records spatial relationships in the scene while efficiently maintaining a dynamic KG.
-Unlike previous KG construction algorithms gutierrez2024hipporag ; edge2024local , which are designed for static information gathering.
-Our spatial memory needs to be updated frequently.
-To speed up the KG update process, we design a two-phase progressive approach: (1) Rapid Response Phase: New information is quickly buffered to avoid losing observations.
-(2) Local Integration Phase: On buffer saturation or conflict detection, affected local subgraphs are integrated (entity de-duplication, relationship merging, conflict detection).
-This algorithm limits KG updates to the environment.
-For each update, we only consider relevant segments, controlling efficiency and updater context length for better performance.
-The algorithm is shown in Algorithm 1.
-Meanwhile, to illustrate the growth of dynamic spatial memory, we provide a sample in Appendix D.
-Our retrieval-based incremental KG update algorithm has provable efficiency guarantees.
-Specifically, for a KG with $n$ nodes and maximum degree $D$, the number of nodes processed in each update is bounded by $O(D^{K})$ where $K$ is the retrieval hop distance (see Appendix C for formal proof).
-This ensures scalability even as the spatial memory grows over time.
-3.2.2 Lifelong Learning System
-In real-world scenarios, agents process sequential tasks over their lifetime, requiring continuous improvement through prior experiences.
-Drawing on cognitive psychology’s classification of human long-term memory, we divide our system into episodic and semantic memory: the former records agent-environment interaction histories, while the latter extracts experiential insights to support long-term task reasoning.
-This update process mirrors human daily experience consolidation during sleep maboudi2024retuning .
+Due to the dynamic and partially observable nature of real embodied environments, completing long-term tasks requires dynamically maintaining a spatial-temporal memory system.
+The system records the agent’s interaction history with the environment and maintains a continuously updated spatial model of the current surroundings.
+To achieve these two purposes, we design spatial memory and temporal memory.
+The combination of spatial and temporal memory provides the embodied agent with comprehensive contextual awareness, supporting effective planning and interactive environmental learning in real-world settings.
+Temporal Memory.
+Interactions between the robot and the environment must be recorded.
+For such temporally sequential memory, a simple structure suffices: a sequential buffer with automatic summarization triggered upon reaching capacity.
+The buffer tracks the high-level actions from the Embodied Agent, how the low-level executor actually controls the robot, and the environmental feedback received.
+In a specific design, temporal memory can store up to $N$ interaction summaries, each generated by an information preprocessor.
+When the buffer is full, rather than discarding old entries outright, we compress the oldest $N$ steps into a single summarized entry using a VLM, which is then reinserted at the front of the buffer, ensuring continuous context retention without unbounded growth.
+Spatial Memory.
+Current spatial memory approaches often rely on RGB-D cameras to reconstruct 3D point clouds (zhang2023building, ; chang2023goat, ).
+However, such representations are overly detailed for high-level planning in embodied agents — precise geometric relationships (e.g., exact distances between objects) are unnecessary.
+Moreover, point clouds are difficult to update efficiently in dynamic scenes.
+To address these problems, we advocate for an abstracted spatial memory that captures object-level entities and their spatial relationships.
+Inspired by gutierrez2024hipporag , we represent this memory as a KG: objects become vertices, and spatial relations are encoded as edges.
+The KG focuses on high-level spatial relations (e.g., “cup on table”, “key left of drawer”).
+Unlike geometric representations that encode precise coordinates, which are often irrelevant for high-level planning, our KG focuses on semantically meaningful, task-relevant relations.
+This abstraction enhances the agent’s spatial reasoning capability in dynamic environments.
+In our memory system, spatial memory is a dynamic KG storing high-level spatial knowledge.
+To overcome the limited spatial reasoning of embodied agents in dynamic environments, we use a KG.
+We introduce a retrieval-driven, incremental KG update algorithm that maintains a locally modifiable, globally consistent, and dynamically adaptive spatial memory.
+As illustrated in the right panel of Figure 2, the update process proceeds in four steps: (1) retrieves the most relevant sub-KG around new observations.
+(2) Injects new relations from the current observation by a VLM-based Relation Retriever.
+(3) Detects and resolves conflicts between newly extracted relations and existing ones (e.g., “cup on table” vs.
+‘cup in drawer”) using a VLM-based resolver, which decides whether to add, delete, or modify edges.
+(4) Merges back and prunes isolated vertices.
+Moreover, our retrieval-based incremental update algorithm comes with provable efficiency guarantees: for a KG with $n$ vertices and maximum degree $D$, the number of vertices processed per update is bounded by $O(D^{K})$, where $K$ is the retrieval hop distance (see Appendix C.3 for formal analysis).
+Further architectural and implementation details are provided in Appendix C.1.
+2.2.2 Interactive Environmental Learning System
+Although episodic memory can record “what happened,” it cannot directly answer “why it succeeded/failed” or “what should be done next time.” Drawing on cognitive psychology’s classification of human long-term memory, we divide our system into episodic and semantic memory: the former records agent-environment interaction histories, while the latter extracts experiential insights to support long-term task reasoning.
+This update process mirrors human daily experience consolidation during sleep (maboudi2024retuning, ).
 Episodic Memory: It captures task-level interactions, accounting for temporal interdependencies between sequential tasks in the same environment.
@@
 Each action is planned and passed on to the low-level executor for execution.
-To enable closed-loop control in embodied environments, the Closed-Loop Planning Module adopts the Planner-Critic mechanism lei2025clea , which consists of the planner and the critic module.
+To enable closed-loop control in embodied environments, the Closed-Loop Planning Module adopts the Planner-Critic mechanism (lei2025clea, ), which consists of the planner and the critic module.
 For each planning step, the planner generates a long-term plan consisting of multiple steps.
@@
 Thus, before executing each step, we use the Critic model to evaluate whether the proposed action in this step remains appropriate under the latest environment.
-If not, the planner will replan based on the latest information.
-The demonstration of this process is shown in Fig.
-2.
+If not, the planner will re-plan based on the latest information.
+The demonstration of this process is shown in Figure 2.
 However, our experiments reveal that the original Planner-Critic mechanism may suffer from infinite loops.
@@
 This ensures that even if the Critic persistently demands replanning, the RoboMemory will still execute actions.
-3.4 Low-level Executer
+2.4 Low-level Executor
 The RoboMemory framework is a two-layer hierarchical agent framework.
@@
 The upper layer is responsible only for high-level planning, while the Low-level Executor carries out the actions planned by the upper layer in the real environment.
-We use $\pi_{0}$ black2024pi_0 as the action executor in the real world: actions planned by RoboMemory are converted into arm and chassis movements via the Low-level Executer, with LoRA fine-tuning applied to optimize performance in real-world robotic tasks.
-More details are provided in Appendix B.
+We employ a LoRA-finetuned VLA model, $\pi_{0}$ (hu2022lora, ; black2024pi_0, ), to generate manipulation actions, and a SLAM-based navigation model for locomotion.
+The low-level executor then translates high-level actions planned by RoboMemory into concrete arm and chassis movements in the real world.
 Table 1: Comparison of Success Rates (SR) and Goal Condition Success Rates (GC) across difficulty levels between RoboMemory and baseline methods on EB-ALFRED.
@@
 InternVL3-78B 37.0% - 38.0% - 36.0% -
-InternVL2.5-38B 23.3% 31.3% 36.0% 37.3% 26.0% 36.5%
+InternVL2.5-38B 31.0% 36.9% 36.0% 37.3% 26.0% 36.5%
 Qwen2.5-VL-72B-Ins 42.0 % - 50.0 % - 34.0 % -
@@
 3 Experiments
-4.1 Experimental Setup
-4.1.1 Benchmarks
-To evaluate the task planning ability of RoboMemory, we select a subset of the EB-ALFRED benchmark from EmbodiedBenchyang2025embodiedbench .
+3.1 Benchmarks
+To evaluate the task planning ability of RoboMemory, we select a subset of the EB-ALFRED benchmark from EmbodiedBench(yang2025embodiedbench, ).
 We selected the Base and Long subsets because they aim to test the agent’s planning ability.
-The Base and Long subset comprises 100 tasks for complex long-term embodied tasks.
-The EB-ALFRED environment provides a visually-grounded operational setting that closely mimics real-world conditions (see appendix B for environment details), enabling direct comparison with established baselines.
-Additionally, we have also evaluated RoboMemory’s capabilities on the EB-Habitat benchmark yang2025embodiedbench . For detailed results, please refer to Appendix A.
-Moreover, we set up a real-world environment to test the lifelong learning ability of RoboMemory in the real world.
-4.1.2 Settings & Baselines
+The Base and Long subset comprises 100 tasks for complex embodied tasks.
+The EB-ALFRED environment provides a visually-grounded operational setting that closely mimics real-world conditions (see Appendix D for environment details), enabling direct comparison with established baselines.
+Additionally, we have also evaluated RoboMemory’s capabilities on the EB-Habitat benchmark (yang2025embodiedbench, ). For detailed results, please refer to Appendix B.
+Moreover, we set up an environment to test the interactive environmental learning ability of RoboMemory in the real world.
+3.2 Settings & Baselines
 To facilitate comparisons, we consider two types of baselines.
@@
 We compare their performance with RoboMemory.
-For closed source VLMs, we choose GPT-4o and GPT-4o-mini GPT-4o-mini ; hurst2024gpt , Claude3.5-Sonnet Claude-3.5-Sonnet , Gemini-1.5-Pro and Gemini-2.0-flash team2024gemini ; Gemini2.0 .
-For open source VLMs, we choose LLama-3.2-90B-Vision-Ins llama3.2 , InternVL-2.5-78B/28B chen2024expanding , InternVL-3-72B zhu2025internvl3 , and bai2025qwen2 .
-Secondly, we choose three agent frameworks: (1) Reflexion shinn2024reflexion , which introduces a simple long-term memory and a self-reflection module.
+For closed source VLMs, we choose GPT-4o and GPT-4o-mini (GPT-4o-mini, ; hurst2024gpt, ), Claude3.5-Sonnet (Claude-3.5-Sonnet, ), Gemini-1.5-Pro and Gemini-2.0-flash (team2024gemini, ; Gemini2.0, ).
+For open source VLMs, we choose Llama-3.2-90B-Vision-Ins (llama3.2, ), InternVL-2.5-78B/28B (chen2024expanding, ), InternVL-3-72B (zhu2025internvl3, ), and Qwen2.5-VL-72B-Ins (bai2025qwen2, ).
+Secondly, we choose three agent frameworks: (1) Reflexion (shinn2024reflexion, ), which introduces a simple long-term memory and a self-reflection module.
 Reflexion uses the self-reflection module to summarize experiences as long-term memory, thereby enhancing the model’s capabilities.
-(2) Voyager wang2023voyager , which utilizes a skill library as its procedural memory, is a widely used baseline for embodied agent planning.
-(3) Cradle tan2024cradle , which proposes a general agent framework with episodic and procedural memory and gains good performances at various multi-model agent tasks.
-In our experiments, each agent framework is tested using Qwen2.5-VL-72b-Ins qwen2.5 .
+(2) Voyager (wang2023voyager, ), which utilizes a skill library as its procedural memory, is a widely used baseline for embodied agent planning.
+(3) Cradle (tan2024cradle, ), which proposes a general agent framework with episodic and procedural memory and gains good performances at various multi-model agent tasks.
+In our experiments, each agent framework is tested using Qwen2.5-VL-72b-Ins (qwen2.5, ).
 The Qwen2.5-VL-72b-Ins represents a high-performing open-source alternative.
-Notably, the Qwen2.5-VL-72b-Ins demonstrates performance comparable to advanced closed-source VLMs in several benchmark tasks white2024livebench .
-We use the Qwen3-Embedding model zhang2025qwen3 to create embedding vectors for RAGs in RoboMemory.
+Notably, the Qwen2.5-VL-72b-Ins demonstrates performance comparable to advanced closed-source VLMs in several benchmark tasks (white2024livebench, ).
+We use the Qwen3-Embedding model (zhang2025qwen3, ) to create embedding vectors for RAGs in RoboMemory.
 For the Low-level Executor, since EB-ALFRED provides high-level action APIs, we use the low-level executor provided by EmbodiedBench instead of the VLA-based method.
-4.1.3 Evaluation metrics
 We define two evaluation metrics to assess the performance: (1) Success Rate (SR), which is the ratio of completed tasks to the total number of tasks in each difficulty level.
@@
 The task is considered successful if $SCN_{x}=GCN_{x}$.
-Figure 3: The reason why RoboMemory failed to complete the task.
+Figure 3: Efficiency improvement of Comprehensive Embodied Memory System
 Table 2: Ablation Study on RoboMemory’s Success Rate (SR)
@@
 - w/o long-term memory 57% 66% 48%
+- w/o episodic memory 62% 68% 56%
+- w/o semantic memory 58% 66% 50%
 3.3 Main results
 As shown in Table 1, our model achieves significant improvements over both single VLM agents and Agent frameworks on the EB-ALFRED.
-Compared to the SOTA Single VLM-Agent model, Claude3.5-Sonnet, RoboMemory with Qwen2.5-VL-72B-Ins backbone improves the average SR by 5% and GC by 15%.
+Compared to the SOTA Single VLM-Agent model, Gemini1.5-Pro, RoboMemory with Qwen2.5-VL-72B-Ins backbone improves the average SR by 3% and GC by 8.7%.
 This demonstrates RoboMemory’s superiority over single VLM-Agents, proving that an Agent framework with open-source models can outperform closed-source SOTA models.
@@
 And it is more robust when encountering unexpected situations.
+3.4 Efficiency Analysis
+To evaluate the efficiency of the Comprehensive Embodied Memory module, we tasked RoboMemory with executing 10 long-horizon tasks, each comprising approximately 50 steps.
+We exclusively measured the wall-clock time consumed by memory update and retrieval operations.
+We analyzed the scaling behavior of memory update latency across three distinct configurations: (1) fully parallel update and retrieval across all memory modules; (2) sequential update of each memory module without parallelism; and (3) update of only the most fundamental memory component: the Temporal Memory.
+Results are presented in Figure 3.
+As shown, our parallel update strategy enables updating a multi-module memory system with latency comparable to that of updating a single base memory module.
+This demonstrates the critical efficiency gains afforded by parallelization across the memory architecture.
 3.5 Ablation Studies
@@
 Results are shown in Table 2.
-4.3.1 Long-term Memory
-Adding long-term memory significantly improved RoboMemory’s success rate.
-The experiment shows that it enables continuous learning while it attempts to complete tasks.
+Long-term Memory Adding long-term memory significantly improved RoboMemory’s success rate.
+The experiment shows that it enables interactive environmental learning while attempting to complete tasks.
 The semantic memory learns low-level skills’ properties, such as in what circumstances an action may fail.
-The temporal memory records all task attempts (successful/failed), providing valuable experience at the task level, giving insight about how to complete a task successfully.
+The temporal memory records all task attempts (successful/failed), providing valuable experience at the task level and giving insight into how to complete a task successfully.
 This helps the RoboMemory predict action outcomes and avoid ineffective attempts.
-This ability indicates that the RoboMemory has lifelong learning capability.
-4.3.2 Spatial Memory
-Spatial memory is crucial for embodied agents, especially given that current pretrained VLMs have limited spatial understanding ability.
+This ability indicates that the RoboMemory has an interactive environmental learning capability.
+Spatial Memory Spatial memory is crucial for embodied agents, especially given that current pretrained VLMs have limited spatial understanding ability.
 Our novel dynamic KG update algorithm enables KG-based spatial memory in dynamic environments.
 This spatial reasoning helps RoboMemory handle partially observable embodied settings.
-4.3.3 Critic Module
-Table 2 shows performance without the critic module (55% vs 67% with full system).
+Critic Module Table 2 shows performance without the critic module (55% vs 67% with full system).
 This drop highlights how the critic’s closed-loop planning adapts to dynamic environments.
 It helps RoboMemory recover from failures faster and handle unexpected situations better.
-4.4 Error Analysis
+Figure 4: Visualization of the experimental environment.
+Figure 5: The improvement of RoboMemory after learning in the real world.
+3.6 Real-world Robot Deployment
+To evaluate RoboMemory’s interactive environmental learning capability in the real world, we designed a kitchen environment inspired by EB-ALFRED and EB-Habitat.
+The scene contains 5 navigable points, 8 interactive objects, and over 10 non-interactive (but potentially distracting) items.
+The environment is shown in Figure 5.
+In the real world, we use interactive environmental video recordings captured during action execution (rather than static snapshots taken after action completion) as RoboMemory’s input.
+This provides a more temporally coherent perception.
+We created three task categories (5 tasks each).
+Tasks are matched to EB-ALFRED’s Base subset (avg.
+oracle: 10–20 steps), though actual executions often exceed 20 steps due to search and error recovery.
+Due to search and error recovery, the robot often exceeds 20 steps per task.
+Additional hardware experiment details are in Appendix D.
+To test the interactive environmental learning ability of RoboMemory, we run each task twice without clearing long-term memory between attempts.
+Success rates for first and second attempts are shown in Figure 5.
+The second attempt showed significantly higher success rates.
+This proves RoboMemory’s long-term memory effectively guides subsequent tasks in real embodied environments.
+Key observations include: (1) Closed-loop error recovery: RoboMemory retries failed actions when possible, even if the low-level executor (VLA model) fails.
+(2) Spatial reasoning: RoboMemory remembers object locations and spatial relationships using its memory.
+(3) Interactive environmental learning: RoboMemory analyzes failure causes reasonably.
+These analyses guide future decisions.
+Detailed examples demonstrating these capabilities and further discussions are provided in Appendix E.
+Moreover, we observed a significant drop in task success rates when deploying the agent with the Low-level Executor in real-world environments.
+This performance degradation primarily stems from the executor’s inherent limitations: (1) The VLA model exhibits unreliable instruction-following capabilities, frequently failing during grasping actions or selecting incorrect objects; (2) Pre-trained VLM models demonstrate inadequate video understanding - while capable of recognizing static objects, they struggle to interpret dynamic visual information such as action failures or state changes.
+These limitations collectively contribute to the reduced performance compared to simulated environments.
+4 Conclusion and Future Work
+In summary, RoboMemory, a brain-inspired multi-memory framework, facilitates long-horizon planning and interactive environmental learning in real-world embodied systems by addressing key challenges such as memory latency, task correlation capture, and planning loops.
+Experiments on EmbodiedBench demonstrate that RoboMemory outperforms state-of-the-art closed-source VLMs and agent frameworks, with ablation studies confirming the critical roles of the Critic module and spatial/long-term memory.
+Real-world deployment further validates its interactive learning capability through improved success rates in repeated tasks.
+Despite limitations arising from reasoning errors and executor dependence, RoboMemory provides a foundation for generalizable, memory-augmented agents, with future work aimed at refining reasoning and enhancing execution robustness.
+A notable open challenge in hierarchical embodied agents, including RoboMemory, lies in the interface between high-level planners and low-level executors.
+Existing frameworks typically rely on language instructions to convey actions, yet some execution details (e.g., precise grasp points) are difficult to describe textually and are better captured through other modalities, such as vision.
+While our current work emphasizes long-term planning and interactive learning, future research may improve generalization by developing richer multimodal interactions between the agent and executor.
+Appendix A Related Work
+A.1 VLM/LLM-based Agentic Frameworks in Embodied Tasks
+The rapid advancement of VLMs/LLMs has led to diverse agent frameworks in embodied environments [42, 33, 24].
+Embodied tasks involve partial observability and long-horizon planning, requiring memory systems to retain context.
+Some use time-ordered context buffers for short-term memory (due to VLMs/LLMs’ limited long-context processing) [42, 29]; others adopt experience buffers as long-term semantic memory [14, 32].
+For long-duration tasks, skill libraries serve as procedural memory, with agents accumulating skills via interaction [38, 35].
+However, in real-world settings, the low-level executor may fail to complete the task, making it challenging to construct a reusable, code-based skill library.
+So, explicit procedural memory still needs to be improved in real-world settings.
+Moreover, Recent efforts integrate diverse memories [44, 35, 1] but focus on virtual/GUI environments, leaving real-world multi-modal memory support for long-term planning under-explored.
+A.2 Vision Language Action Model
+Current work on VLA models uses imitation learning to output low-level controls from language and visuals [5, 47, 4, 22] but is limited to tabletop tasks and single actions, restricting long-horizon planning.
+VLAs lack long-term execution abilities, while high-level agents excel at planning.
+Recent works combine high-level frameworks with VLA executors, some augmented with simple memory [31, 34, 43, 41] for longer tasks.
+However, real-world robots need more sophisticated memory to handle continuous multi-task operations over extended periods.
+A.3 Memory Frameworks
+Many previous works improve long-term planning via memory systems: Voyager [38] uses a skill library in Minecraft but lacks diverse memory types; CoELA [44] includes procedural, semantic, and episodic memory with a task-specific 2D map; MSI-Agent [14] utilizes insight as long-term memory for in-task learning.
+Hippo Retrieval Augmented Generation (RAG) [18] mimics the hippocampus and introduces KGs as long-term memory indices [6, 9], enhancing retrieval.
+However, the previous approach is mainly focused on constructing a KG with a static long context, such as a book, but it is hard to update the graph.
+We need to update the information in KG for the embodied task.
+Our approach builds a more general LLM-based memory system using a dynamic KG like Hippo RAG, which is designed for embodied tasks.
+Furthermore, we summarize the differences among different memory systems in previous work.
+The comparison is shown in the Table 3.
+Table 3: Comparison of Memory-Related Methods in Embodied Agent
+Method Multimodal Episodic Semantic Spatial Temporal Procedural Memory Implementation Real Robot
+NeSyC [12] Symbolic logic rules
+Reflexion [32] Buffer
+Voyager [38] RAG
+MSI-Agent [14] Database, RAG
+CoELA [44] Top-down semantic map
+Cradle [35] RAG
+Agent-S [1] RAG
+Expel [46] Buffer
+AutoManual [8] Buffer
+HiRobot [31] /
+Being-0 [43] Buffer
+RoboOS [34] Scene graph, database
+RoboMemory (Ours) RAG,KG
+Appendix B Additional Experiments
+B.1 Error Analysis
 We summarize the common errors of RoboMemory in the previous experiments. We classify errors into three main types: planning errors, reasoning errors, and perception errors.
@@
 Perception errors occur when incorrect information is provided to the planner-critic module.
+Figure 6: The reason why RoboMemory failed to complete the task.
 We analyze RoboMemory trajectories for failed tasks.
@@
 We calculate the occurrence probability of each error type to show RoboMemory’s strengths and weaknesses.
-The results are shown in Fig.
-3.
+The results are shown in Figure 6.
 We can observe that among all error types, the planning errors are the most common.
@@
 We can observe that although some hallucinations can be handled by the critic module or memory information, there are still some cases in which the planner ignores all insights from memory and critic and fails to complete the task.
-The detailed examples and discussions are provided in Appendix D.
-Figure 4: Visualization of the experimental environment.
-Figure 5: The improvement of RoboMemory after learning in the real world.
-4.5 Real-world Robot Deployment
-4.5.1 Environmental Setup
-To evaluate RoboMemory’s lifelong learning capability in the real world, we designed a kitchen environment inspired by EB-ALFRED and EB-Habitat.
-The scene contains 5 navigable points, 8 interactive objects, and over 10 non-interactive (but potentially distracting) items.
-The environment is shown in Fig.
-5.
-In the real world, we use continuous video recordings during action execution (instead of snapshots after the action execution) as RoboMemory’s input.
-This provides a more temporally coherent perception.
-We created three task categories (5 tasks each, totaling 15 tasks) with difficulty matching EB-ALFRED’s Base subset (average oracle trajectory length from 5 to 10).
-Additional hardware experiment details are in Appendix B.
-To test the lifelong learning ability of RoboMemory, we run each task twice without clearing long-term memory between attempts. Success rates for first and second attempts are shown in Fig. 5.
-4.5.2 Main results
-The second attempt showed significantly higher success rates. This proves RoboMemory’s long-term memory effectively guides subsequent tasks in real embodied environments. Key observations include:
-  1. 1.
-    Closed-loop error recovery: RoboMemory retries failed actions when possible, even if the low-level executor (VLA model) fails.
-  2. 2.
-    Spatial reasoning: RoboMemory remembers object locations and spatial relationships using its memory.
-  3. 3.
-    Lifelong learning: RoboMemory analyzes failure causes reasonably. These analyses guide future decisions.
-Detailed examples demonstrating these capabilities and further discussions are provided in Appendix D.
-Moreover, we observed a significant drop in task success rates when deploying the agent with the Low-level Executor in real-world environments.
-This performance degradation primarily stems from the executor’s inherent limitations: (1) The VLA model exhibits unreliable instruction-following capabilities, frequently failing during grasping actions or selecting incorrect objects; (2) Pre-trained VLM models demonstrate inadequate video understanding - while capable of recognizing static objects, they struggle to interpret dynamic visual information such as action failures or state changes.
-These limitations collectively contribute to the reduced performance compared to simulated environments.
-5 Conclusion and Future Work
-In conclusion, RoboMemory, a brain-inspired multi-memory framework, enables lifelong learning and long-term planning in real-world embodied systems by addressing key challenges: continuous learning, memory latency, task correlation capture, and planning infinite loops.
-Experimental results on EmbodiedBench show it outperforms SOTA closed-source VLMs and agent frameworks, with ablation studies validating critical components like the Critic module and spatial/long-term memory.
-Real-world deployment confirms its lifelong learning capability via improved success rates in repeated tasks.
-While limited by reasoning errors and executor reliance, it paves the way for generalizable memory-augmented agents, with future work focusing on refining reasoning and enhancing execution robustness.
-A key unsolved problem in current hierarchical agent research for embodied tasks, including ours, is about the interaction between high-level agents and low-level executors (e.g., VLAs).
-Most existing frameworks use language instructions solely as action instructions from high-level agents.
-However, some action details are hard to describe with language.
-Other modalities (e.g., vision) can better represent these details (e.g., grasp points).
-While our work focuses on long-term planning and lifelong learning for agents, future work can consider enhancing generalization within existing frameworks by improving the interaction method between the VLA and the Agent.
-Appendix A Additional Experiments
-Figure 6: Comparison of Success Rates (SR) and Goal Condition Success Rates (GC) across difficulty levels between RoboBrain and baseline methods on EB-Habitat.
-A.1 Evaluate on EB-Habitat
+The detailed examples and discussions are provided in Appendix E.
+B.2 Evaluate on EB-Habitat
+Figure 7: Comparison of Success Rates (SR) and Goal Condition Success Rates (GC) across difficulty levels between RoboMemory and baseline methods on EB-Habitat.
 Similar to EB-ALFRED, we also deploy RoboMemory to EB-Habitat.
@@
 Each subset contains 50 different trails.
-The results are shown in Fig 6.
+The results are shown in Figure 7.
 The results demonstrate that our RoboMemory can adapt well to different environments.
@@
 These improvements indicate that RoboMemory enhances the agent’s embodied intelligence in various environments. The key factor is its complete memory system.
+B.3 Additional Efficiency analysis
+We analyze the evolution of the spatial KG during long trajectories in EB-ALFRED, focusing on the first 20 iterations (with 95% confidence intervals).
+As shown in Figure 8, the total number of spatial relationships in the KG (red line) increases gradually over iterations as RoboMemory is exploring the environment.
+In contrast, the number of relationships retrieved for update at each iteration (blue line) remains relatively stable, typically ranging around 10 edges per iteration.
+This stability is achieved because our method only updates a local subgraph relevant to the current observation.
+We define the retrieval ratio as the proportion of relationships updated at each iteration relative to the total number of relationships in the KG.
+As shown in Figure 8, this ratio (illustrated by gray bars) decreases steadily from 76% initially to 28% at iteration 20.
+This trend indicates that, as the KG grows, each update affects a progressively smaller fraction of the entire graph.
+This demonstrates that our spatial KG update mechanism effectively localizes modifications, ensuring computational efficiency and mitigating interference through context-aware incremental updates.
+Figure 8: Average relationships related to update in spatial memory in each step.
+Appendix C Dynamic Spatial Memory Update Algorithm
+C.1 Detailed Algorithm of Spatial KG Update
+Algorithm 1 Retrieval-based Incremental Knowledge Graph Update Algorithm
+0: New spatial knowledge graph $G_{\text{new}}=(V_{\text{new}},E_{\text{new}})$, main spatial knowledge graph $G=(V,E)$, queries $q\in Q$, entity & query embeddings $\mathcal{E}:V\cup Q\rightarrow\mathbb{R}^{d}$, maximum number of retrieved vertices $n$, maximum k hops $k$, vlm-base conflict resolver $\text{ResolveConflict}(\cdot)$
+0: Updated consistent knowledge graph $G^{\prime}$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+9: Remove isolated vertices from $G^{\prime}$
+10:return $G^{\prime}$
+Spatial Memory is a dynamically updated KG-based module designed to overcome agents’ limitations in spatial reasoning.
+Specifically, our Spatial Memory is formulated as a directed KG $G=(V,E)$, where $V$ denotes the set of all objects in the environment.
+Each object is a vertex in the KG.
+Each object’s name is encoded into a single semantic embedding vector via a pretrained embedding model [45].
+The edge set $E$ captures spatial relationships between objects, each represented as a triple (e.g., $obj_{1},\text{relationship},obj_{2}]$).
+To update the Spatial KG, we need to continuously extract new relationships from current observations and update the old relationships in the Spatial KG.
+We use a VLM-based conflict resolver to address this problem.
+However, the more relationships provided to the conflict resolver, the more time it needs to update the KG.
+So we need to update relationships that are only related to the current situation.
+We design an algorithm that retrieves a sub-graph of KG that includes all vertices related to the current situation and both old and new relationships among them.
+We provide the sub-graph and new relationships to the conflict resolver.
+We need the conflict resolver to update the sub-graph based on information from the new relationships.
+The algorithm is shown in Algorithm 1.
+To update $G$, we make use of the information provided by the step summarizer and query generator from the information preprocessor introduced in Sec.
+2.1.
+First, a pretrained VLM-based Relation Retriever extracts the latest spatial relationships $G_{\text{new}}=(V_{\text{new}},E_{\text{new}})$ from the information provided by the step summarizer, which records high-level information in the current observation.
+Next, natural language queries (represented as $q\in Q$) (provided by the query generator) are used to retrieve relevant object vertices from $G$ via cosine similarity search.
+We select the top $n$ similar vertices compared with the query.
+These vertices are represent by $V_{\text{similar}}$.
+Spatial KG maintains the relationships among different objects, so if we want to retrieve spatial information from spatial KG, we need to search for other objects that are related to the objects we observed in the current observation.
+In this way, we not only remember objects we can see, but also know the spatial information of the objects we cannot see.
+So we choose the k-hop algorithm to expand $V_{\text{similar}}$ using a K-hop neighborhood algorithm to capture contextually related objects.
+The K-hop algorithm is represented as $\text{K-hop}_{k}(V,G)$, which returns all vertices reachable within $\leq k$ hops from any vertices in $V$.
+The retrieved vertices are $V_{\text{expand}}$.
+We combine the vertices retrieved by cosine similarity and their K-hop neighbors to $V_{\text{retrieved}}$.
+However, we need to resolve the conflict between new and old relationships.
+So $V_{\text{retrieved}}$ and the relationships among $V_{\text{retrieved}}$ is not enough.
+We need new vertices and relationships involved in the graph we provided to the VLM-based conflict resolver.
+To extract all vertices and relationships for the VLM-based conflict resolver and relationships, we not only need the relationships from KG (old information) and $G_{\text{new}}$, which represent new information.
+We need to connect old information and new information.
+To achieve this goal, we merge $G_{\text{new}}$ to $G$, which add new edges and vertices to $G$.
+We denote the merged KG as $G_{\text{union}}$.
+In $G_{\text{union}}$, we mix out-of-date and latest information.
+Then, we extract an induced sub-graph of $V_{\text{merged}}=V_{\text{retrieved}}\cup V_{\text{new}}$ from $G_{\text{union}}$.
+As both vertices from old graph $G$ and new graph is mixed in $V_{\text{merged}}$ and both edges from old and new graph is in $G_{\text{union}}$, the retrieved induced sub-graph $G_{\text{local}}$ contains all out-of-date and latest relationships among vertices that is related to current situation.
+As $G_{\text{local}}$ contains all out-of-date and latest relationships and those relationships may have conflicts, a VLM-based conflict resolver (represent as $\text{ResolveConflict}(\cdot)$) is designed to resolve conflicts in $G_{\text{local}}$, and make sure that the relationship is the latest.
+The conflict resolver will take in $G_{\text{new}}$ and $G_{\text{local}}$, where $G_{\text{local}}$ is the graph waiting for update and $G_{\text{new}}$ provide update signal.
+The VLM-based conflict resolver will perform necessary updates such as adding vertices or inserting, deleting, or modifying relationships in $G_{\text{local}}$ based on $G_{\text{new}}$.
+The reconciled subgraph is then merged back into $G$, and any vertices that have lost all connections to other vertices during the update are pruned.
+This design offers two key advantages: (1) Efficiency via localized updates: By restricting modifications to a context-relevant subgraph, we significantly reduce the number of relationships processed per update.
+Since VLMs struggle with reasoning over large sets of relationships, this constraint substantially improves both the efficiency and effectiveness of VLM-based KG updates.
+(2) Dynamic adaptability: The system continuously maintains up-to-date spatial knowledge, enabling agents to operate robustly in dynamic real-world environments.
+Figure 9: Visualization of Spatial Memory’s dynamic update process.
+C.2 Example of Dynamic Spatial Memory Update process
+In RoboMemory’s Spatial Memory, the KG is dynamically constructed during environment exploration.
+As illustrated in Figure 9, we demonstrate the progressive expansion of the KG in Spatial Memory as the agent navigates through the environment.
+The figure indicates a continuous growth in the number of both vertices and edges of the KG as exploration progresses.
+Notably, the KG undergoes dynamic updates through RoboMemory’s environmental interactions.
+For example, the initial KG state displays the relation “I am near the apple.
+But as the agent picks up the apple in the third step, in the fourth KG, the relationship becomes “I hold the apple”.
+This demonstrates RoboMemory’s capability for dynamic KG maintenance and expansion.
+By querying this KG, the Planner-Critic module gains access to rich spatial information, empowering RoboMemory with robust spatial memory capabilities that significantly enhance its performance in both TextWorld and EmbodiedBench environments.
+C.3 Proof of Dynamic Spatial Memory Update Algorithm
+Theorem 1 (Upper Bound on K-hop Vertex Extraction in Directed Graphs).
+Let $G=(V,E)$ be a finite directed graph with maximum out-degree $D\geq 1$, and let $\mathcal{S}\subseteq V$ be a set of $M$ source vertices.
+Define the K-hop neighborhood $\mathcal{N}_{K}(s)$ of a vertex $s\in\mathcal{S}$ as the set of vertices reachable from $s$ via directed paths of length at most $K$.
+Then the total number of distinct vertices in the union of all K-hop neighborhoods,
+$\mathcal{N}_{K}(\mathcal{S})=\bigcup_{s\in\mathcal{S}}\mathcal{N}_{K}(s),$
+Satisfies the following upper bound:
+$|\mathcal{N}_{K}(\mathcal{S})|\leq\begin{cases}M\cdot\dfrac{D^{K+1}-1}{D-1},&\text{if }D>1,\\ M\cdot(K+1),&\text{if }D=1.\end{cases}$
+Proof.
+For any vertex $s\in\mathcal{S}$, the number of distinct vertices reachable from $s$ within $i$ hops is at most $D^{i}$, assuming the worst-case scenario where each vertex encountered has the maximum out-degree $D$, and all neighbors are distinct and non-overlapping.
+Thus, the size of the K-hop neighborhood of a single vertex satisfies:
+$|\mathcal{N}_{K}(s)|\leq\sum_{i=0}^{K}D^{i}=\begin{cases}\dfrac{D^{K+1}-1}{D-1},&\text{if }D>1,\\ K+1,&\text{if }D=1.\end{cases}$
+Since there are $M$ such source vertices and assuming no overlaps between their K-hop neighborhoods (worst case), the union size satisfies:
+$|\mathcal{N}_{K}(\mathcal{S})|\leq M\cdot|\mathcal{N}_{K}(s)|.$
+Substituting the bound on $|\mathcal{N}_{K}(s)|$ gives the result. ∎
+Theorem 2 (Upper Bound for K-hop Vertex Extraction in Normalized Directed Graphs).
+Let $G=(V,E)$ be a finite directed graph with $|V|=n$ vertices.
+Assume the maximum out-degree is at most $D_{\max}=Dn$, and the maximum in-degree is at most $N_{\max}=Nn$, where $D,N\in(0,1]$ are constants.
+Let $\mathcal{S}\subseteq V$ be a set of $M$ source vertices.
+Define $\mathcal{N}_{K}(\mathcal{S})$ as the union of all vertices reachable from $\mathcal{S}$ via paths of length at most $K$, using only outgoing edges.
+Then the number of extracted vertices satisfies:
+$|\mathcal{N}_{K}(\mathcal{S})|\leq\min\left\{n,\;M\cdot\frac{(Dn)^{K+1}-1}{Dn-1}\right\}.$
+In particular, when $Dn\gg 1$, we have the approximation:
+$|\mathcal{N}_{K}(\mathcal{S})|\lessapprox M\cdot(Dn)^{K}.$
+Proof.
+For each vertex $s\in\mathcal{S}$, the maximum number of reachable vertices within $i$-hops is at most $(Dn)^{i}$ under the assumption of maximum out-degree and no overlap.
+Summing over hops from 0 to $K$, we get for each root:
+$|\mathcal{N}_{K}(s)|\leq\sum_{i=0}^{K}(Dn)^{i}=\frac{(Dn)^{K+1}-1}{Dn-1}.$
+Assuming no overlap among the $M$ source vertex expansions (worst case), we have:
+$|\mathcal{N}_{K}(\mathcal{S})|\leq M\cdot\frac{(Dn)^{K+1}-1}{Dn-1}.$
+Since the total number of vertices in the graph is $n$, this quantity is also trivially bounded above by $n$, yielding the result. ∎
+Appendix D Additional Environment settings
+D.1 EB-ALFRED and EB-Habitat
+We adopt the same environment parameters as in EmbodiedBench.
+The maximum steps per task are set to 30, with image inputs of size 500 $\times$ 500.
+The temporal memory buffer length is set to 3.
+However, we modified the action formats of EB-ALFRED and EB-Habitat to simulate real-world scenarios better.
+Specifically, we define different action APIs (Python functions), where each action takes an object parameter indicating its target.
+We extract all possible objects from the environment as inputs to the Agent.
+The Agent must select appropriate actions and object parameters based on task requirements.
+Compared to the original interaction method in EmbodiedBench (which enumerates all possible actions, including both action names and target objects, and requires the Agent to choose), our approach offers greater flexibility.
+The detailed action APIs are presented in Table 4.
+Since EB-ALFRED and EB-Habitat provide comprehensive high-level action APIs, we do not employ the VLA-Based Low-Level Executor in these environments.
+Instead, we utilize the built-in low-level controllers from EmbodiedBench.
+D.2 Real-world experiments
 Table 4: Robot Action Command For different environments
 Action Type EB-ALFRED EB-Habitat Real World
-Navigate to object/navigation point find(obj) navigate(point) navigate_to(point)
+Navigation find(obj) navigate(point) navigate_to(point)
 Pick Up Object pick_up(obj) pick(obj) pick_up(obj)
@@
 Task Complete task_complete()
-Appendix B Additional Environment settings
-B.1 EB-ALFRED and EB-Habitat
-We adopt the same environment parameters as in EmbodiedBench.
-The maximum steps per task are set to 30, with image inputs of size 500 $\times$ 500.
-The temporal memory buffer length is set to 3.
-However, we modify the action formats of EB-ALFRED and EB-Habitat to better simulate real-world scenarios.
-Specifically, we define different action APIs (Python functions), where each action takes an object parameter indicating its target.
-We extract all possible objects from the environment as inputs to the Agent.
-The Agent must select appropriate actions and object parameters based on task requirements.
-Compared to the original interaction method in EmbodiedBench (which enumerates all possible actions, including both action names and target objects, and requires the Agent to choose), our approach offers greater flexibility.
-The detailed action APIs are presented in Table 3.
-Since EB-ALFRED and EB-Habitat provide comprehensive high-level action APIs, we do not employ the VLA-Based Low-Level Executor in these environments.
-Instead, we utilize the built-in low-level controllers from EmbodiedBench.
-B.2 Real-world experiments
-We construct a common kitchen scenario to evaluate the RoboMemory framework’s lifelong learning capabilities in real-world settings.
+We construct a common kitchen scenario to evaluate the RoboMemory framework’s interactive environmental learning capabilities in real-world settings.
 Using Mobile ALOHA [15] as our physical robotic platform, we design three categories of tasks: (1) Pick up & put down: The agent must locate a specified object among all possible positions and place it at a designated location.
@@
 This task evaluates the agent’s understanding of object relationships, requiring it to remember the positions of at least two objects (the container and the target item) and their spatial relationship.
+For each type of task, we design 5 tasks.
+So our experiments include 15 long-term real-world tasks.
 To adapt to the real-world setup, we define high-level action APIs similar to those in EB-ALFRED and EB-Habitat.
@@
 For the low-level executor, we use one main camera and two arm-mounted cameras as input, each with a resolution of 640 $\times$ 480. The temporal memory buffer length is set to 3.
-In our experiments, we set the maximum steps per task to 15.
+In our experiments, we set the maximum steps per task to 25.
 We also provide an API for actively terminating tasks.
 Since real-world environments lack direct success/failure feedback, RoboMemory must autonomously determine task completion.
-To prevent excessively long task execution, we enforce termination after 15 steps if no success is achieved.
+To prevent excessively long task execution, we enforce termination after 25 steps if no success is achieved.
 A single main camera (640 $\times$ 480 resolution) records video during action execution as input for RoboMemory’s higher-level processing.
@@
 Total episodes 1040 Training time 12 hours
-B.3 Training Details of Low-Level Executer
+D.3 Training Details of Low-Level Executor
 We use the $\pi_{0}$ model as our foundation model.
 We collected 1,040 data samples over 10 types of tasks for fine-tuning.
-We use LoRA [17] fine-tuning to save resources during fine-tuning.
+We use LoRA fine-tuning to save resources during fine-tuning.
 The specific fine-tuning parameters and action types are given in Table 5.
@@
 We used SLAM to navigate between these navigation points.
-Appendix C Proof of Dynamic Spatial Memory Update Algorithm
-Theorem 1 (Upper Bound on K-hop Node Extraction in Directed Graphs).
-Let $G=(V,E)$ be a finite directed graph with maximum out-degree $D\geq 1$, and let $\mathcal{S}\subseteq V$ be a set of $M$ source nodes.
-Define the K-hop neighborhood $\mathcal{N}_{K}(s)$ of a node $s\in\mathcal{S}$ as the set of nodes reachable from $s$ via directed paths of length at most $K$.
-Then the total number of distinct nodes in the union of all K-hop neighborhoods,
-$\mathcal{N}_{K}(\mathcal{S})=\bigcup_{s\in\mathcal{S}}\mathcal{N}_{K}(s),$
-Satisfies the following upper bound:
-$|\mathcal{N}_{K}(\mathcal{S})|\leq\begin{cases}M\cdot\dfrac{D^{K+1}-1}{D-1},&\text{if }D>1,\\ M\cdot(K+1),&\text{if }D=1.\end{cases}$
-Proof.
-For any node $s\in\mathcal{S}$, the number of distinct nodes reachable from $s$ within $i$ hops is at most $D^{i}$, assuming the worst-case scenario where each node encountered has the maximum out-degree $D$, and all neighbors are distinct and non-overlapping.
-Thus, the size of the K-hop neighborhood of a single node satisfies:
-$|\mathcal{N}_{K}(s)|\leq\sum_{i=0}^{K}D^{i}=\begin{cases}\dfrac{D^{K+1}-1}{D-1},&\text{if }D>1,\\ K+1,&\text{if }D=1.\end{cases}$
-Since there are $M$ such source nodes and assuming no overlaps between their K-hop neighborhoods (worst case), the union size satisfies:
-$|\mathcal{N}_{K}(\mathcal{S})|\leq M\cdot|\mathcal{N}_{K}(s)|.$
-Substituting the bound on $|\mathcal{N}_{K}(s)|$ gives the result. ∎
-Theorem 2 (Upper Bound for K-hop Node Extraction in Normalized Directed Graphs).
-Let $G=(V,E)$ be a finite directed graph with $|V|=n$ nodes.
-Assume the maximum out-degree is at most $D_{\max}=Dn$, and the maximum in-degree is at most $N_{\max}=Nn$, where $D,N\in(0,1]$ are constants.
-Let $\mathcal{S}\subseteq V$ be a set of $M$ source nodes.
-Define $\mathcal{N}_{K}(\mathcal{S})$ as the union of all nodes reachable from $\mathcal{S}$ via paths of length at most $K$, using only outgoing edges.
-Then the number of extracted nodes satisfies:
-$|\mathcal{N}_{K}(\mathcal{S})|\leq\min\left\{n,\;M\cdot\frac{(Dn)^{K+1}-1}{Dn-1}\right\}.$
-In particular, when $Dn\gg 1$, we have the approximation:
-$|\mathcal{N}_{K}(\mathcal{S})|\lessapprox M\cdot(Dn)^{K}.$
-Proof.
-For each node $s\in\mathcal{S}$, the maximum number of reachable nodes within $i$-hops is at most $(Dn)^{i}$ under the assumption of maximum out-degree and no overlap.
-Summing over hops from 0 to $K$, we get for each root:
-$|\mathcal{N}_{K}(s)|\leq\sum_{i=0}^{K}(Dn)^{i}=\frac{(Dn)^{K+1}-1}{Dn-1}.$
-Assuming no overlap among the $M$ source node expansions (worst case), we have:
-$|\mathcal{N}_{K}(\mathcal{S})|\leq M\cdot\frac{(Dn)^{K+1}-1}{Dn-1}.$
-Since the total number of nodes in the graph is $n$, this quantity is also trivially bounded above by $n$, yielding the result. ∎
-Appendix D Supplementary Examples for Qualitative Analysis
-Figure 7: Visualization of Spatial Memory’s dynamic update process.
-D.1 Example of Dynamic Spatial Memory Update process
-In RoboMemory’s Spatial Memory, the KG is dynamically constructed during environment exploration.
-As illustrated in Fig.
-7, we demonstrate the progressive expansion of the KG in Spatial Memory as the agent navigates through the environment.
-The figure indicates a continuous growth in the number of both nodes and edges of the KG as exploration progresses.
-Notably, the KG undergoes dynamic updates through RoboMemory’s environmental interactions.
-For example, the initial KG state displays the relation “I am near the apple.
-But as the agent picks up the apple in the third step, in the fourth KG, the relationship becomes “I hold the apple”.
-This demonstrates RoboMemory’s capability for dynamic KG maintenance and expansion.
-By querying this KG, the Planner-Critic module gains access to rich spatial information, empowering RoboMemory with robust spatial memory capabilities that significantly enhance its performance in both TextWorld and EmbodiedBench environments.
-D.2 Real World
-Figure 8: Case that a task is failed but the experience can help RoboMemory to succeed in the next try.
-In Fig.
-8, we demonstrate an example of RoboMemory learning through trial and error in a real-world environment.
+D.4 Hyperparameters of RoboMemory
+In this section, we describe the hyperparameter settings for the upper brain of RoboMemory: the information preprocessor and the Comprehensive Embodied Memory.
+Importantly, we use a unified set of hyperparameters across all experimental settings, including EB-ALFRED, EB-Habitat, and real-world deployments.
+The information preprocessor consists of two components: a step summarizer and a query generator.
+Given multimodal inputs at each step, the step summarizer produces a single natural language description, while the query generator concurrently formulates 454-5 distinct natural language queries.
+The Comprehensive Embodied Memory integrates four memory modules: Temporal, Spatial, Semantic, and Episodic Memory.
+The Temporal Memory is implemented as a fixed-size buffer with a maximum capacity of $4$ entries.
+For Spatial Memory, during similarity-based retrieval, we first identify the top $N=3$ most relevant vertices and then perform a K-hop graph traversal with $K=2$.
+The Episodic Memory retrieves the top $N=5$ most relevant past experiences for each query.
+The Semantic Memory maintains hierarchical summaries at both the action and task levels; during retrieval, it returns $N_{s}=2$ action-level and $N_{t}=2$ task-level summaries.
+Furthermore, memory updates (e.g., insertion, modification, or deletion) are applied only to the top $N_{\text{update}}=10$ most relevant entries in the Semantic Memory to ensure efficiency and coherence.
+Appendix E Supplementary Examples for Qualitative Analysis
+E.1 Real World
+Figure 10: Case that a task is failed, but the experience can help RoboMemory to succeed in the next try.
+In Figure 10, we demonstrate an example of RoboMemory learning through trial and error in a real-world environment.
 Our task is “place a banana into the oven.” This task required RoboMemory to complete the objectives of finding the banana, picking it up, and transporting it to the oven.
 We observed that RoboMemory became stuck in an infinite loop during the first attempt.
-The banana was randomly placed on the “kitchen counter,” but RoboMemory overlooked this navigation target and remained trapped exploring other navigation targets instead.
+The banana was randomly placed on the “kitchen counter,” but RoboMemory overlooked this navigation target and remained trapped, exploring other navigation targets instead.
 However, based on this bad attempt, the semantic memory summarized that the robot should not repeatedly search in locations where the “banana” could not be found.
@@
 We also provide an example that completes the task in the first attempt.
-The example is shown in Fig.
-9.
+The example is shown in Figure 11.
 This example demonstrates that the RoboMemory has the ability to handle some relatively complex tasks in the real world.
@@
 With the help of the spatial memory, RoboMemory completes the task successfully.
-D.3 EB-ALFRED
+E.2 EB-ALFRED
 We select three examples in EB-ALFRED to show the errors that RoboMemory may encounter and the reasons why or why not RoboMemory can achieve the goal.
-D.3.1 Successful example
+E.2.1 Successful example
 Figure 12: Case that a task is successful with the help of the critic and spatial memory modules.
-We select a successful example to show how RoboMemory performed in the EB-ALFRED environment. The example trajectory is shown in Fig. 10.
+We select a successful example to show how RoboMemory performed in the EB-ALFRED environment. The example trajectory is shown in Figure 12.
 The task of this example is “set a plate with a spoon on it on the kitchen table”.
@@
 In this example, RoboMemory successfully overcame the hallucination and eventually achieved the goal. This example demonstrates that the critic module can help RoboMemory to overcome error cases.
-D.3.2 Failed example
+E.2.2 Failed example
 Figure 13: Case that a task fails in an infinite loop because the critic module failed to stop the agent when its planned action is no longer suitable.
 We demonstrate a representative example of the Critical Error.
-The example trajectory is shown in Fig.
-11.
+The example trajectory is shown in Figure 13.
 In this example, the task involves slicing and heating a tomato and moving the heated tomato slice to the trash can.
@@
 Besides, we provide another example demonstrating a representative failure caused by inaccurate action planning.
-The example trajectory is shown in Fig.
-12.
+The example trajectory is shown in Figure 14.
 In the trajectory, RoboMemory is asked to place two CDs into the drawer.

### v6

@@
 Abstract
-Embodied agents face persistent challenges in real-world environments, including partial observability, limited spatial reasoning, and high-latency multi-memory integration.
-We present RoboMemory, a brain-inspired framework that unifies Spatial, Temporal, Episodic, and Semantic memory under a parallelized architecture for efficient long-horizon planning and interactive environmental learning.
-A dynamic spatial knowledge graph (KG) ensures scalable and consistent memory updates, while a closed-loop planner with a critic module supports adaptive decision-making in dynamic settings.
-Experiments on EmbodiedBench show that RoboMemory, built on Qwen2.5-VL-72B-Ins, improves average success rates by 25% over its baseline and exceeds the closed-source state-of-the-art (SOTA) Gemini-1.5-Pro by 3%.
-Real-world trials further confirm its capacity for cumulative learning, with performance improving across repeated tasks.
-These results highlight RoboMemory as a scalable foundation for memory-augmented embodied intelligence, bridging the gap between cognitive neuroscience and robotic autonomy.
+Embodied intelligence aims to enable robots to learn, reason, and generalize robustly across complex real-world environments.
+However, existing approaches often struggle with partial observability, fragmented spatial reasoning, and inefficient integration of heterogeneous memories, limiting their capacity for long-horizon adaptation.
+To address this, we introduce RoboMemory, a brain-inspired framework that unifies Spatial, Temporal, Episodic, and Semantic memory within a parallelized architecture for efficient long-horizon planning and interactive learning.
+Its core innovations are a dynamic spatial knowledge graph for scalable, consistent memory updates and a closed-loop planner with a critic module for adaptive decision-making.
+Extensive experiments on EmbodiedBench show that RoboMemory, instantiated with Qwen2.5-VL-72B-Ins, improves the average success rate by 26.5% over its strong baseline and even surpasses the closed-source SOTA, Claude-3.5-Sonnet.
+Real-world trials further confirm its capability for cumulative learning, with performance consistently improving over repeated tasks.
+Our results position RoboMemory as a scalable foundation for memory-augmented embodied agents, bridging insights from cognitive neuroscience with practical robotic autonomy.
 Figure 1: RoboMemory adopts a brain-inspired architecture that maps neural components to agent modules, enabling long-term planning and interactive learning across diverse environments (real-world, Habitat, ALFRED) and robotic hardware.
-1 Introduction
-Recent advances in Vision-Language Models (VLMs) (hurst2024gpt, ; bai2025qwen2, ) have enabled their growing use in embodied tasks (park2023generative, ; hu2023toward, ).
+**footnotetext: Equal contribution
+I Introduction
+Recent advances in Vision-Language Models (VLMs) [23, 6] have enabled their growing use in embodied tasks [33, 22].
 VLM-based embodied agents can process multimodal inputs and generate high-level textual commands (e.g., “Pick up the cup”), which require translation via tool APIs to become executable robot actions.
-In contrast, Vision-Language-Action models (VLAs) (kim2024openvla, ; black2024pi_0, ; bjorck2025gr00t, ; chi2023diffusion, ) produce low-level control signals directly but generally rely only on the latest observation.
+In contrast, Vision-Language-Action models (VLAs) [24, 8, 7, 13] produce low-level control signals directly but generally rely only on the latest observation.
 This limits their ability to perform long-horizon, multi-step tasks that require reasoning over task history.
@@
 This highlights a key gap inherent in two distinct technical approaches to embodied intelligence.
-To bridge this gap, recent work (yuan2025being, ; shi2025hi, ; tan2025roboos, ) proposes a “VLM planner + VLA executor” paradigm.
+To bridge this gap, recent work [50, 34, 38] proposes a “VLM planner + VLA executor” paradigm.
 Here, VLM-based embodied agents serve as high-level planners that decompose complex tasks (e.g., “make a coffee”) into executable sub-instructions (e.g., “grasp the cup”) that VLAs can complete.
@@
 This necessitates a planner with robust spatial awareness and long-term memory to maintain a consistent spatial awareness across viewpoints.
-However, most VLM-based agents rely on chat-style context windows (e.g., logging instruction – feedback pairs (yao2022react, )), which lack mechanisms for maintaining an overview of the environment’s spatial layout.
+However, most VLM-based agents rely on chat-style context windows (e.g., logging instruction – feedback pairs [49]), which lack mechanisms for maintaining an overview of the environment’s spatial layout.
 Consequently, agents cannot reliably track object locations or recognize previously visited states.
 Second, pretrained VLMs are rarely trained on embodied planning trajectories, especially long-horizon, spatially grounded ones.
-So VLM-based agents often struggle to generalize to real-world settings (yang2025embodiedbench, ).
+So VLM-based agents often struggle to generalize to real-world settings [47].
 To overcome these challenges, VLM-based planners must support interactive environmental learning — the ability to acquire, integrate, and retrieve spatial, episodic, and semantic knowledge during task execution, thereby enabling adaptation through experience.
-To support long-horizon planning and interactive learning in real-world settings, agents require a comprehensive memory system with multiple specialized modules.
-Recent frameworks (tan2024cradle, ; glocker2025llm, ; wang2023voyager, ; agashe2024agent, ; fu2024msi, ; zhao2024expel, ; chen2024automanual, ) have integrated Retrieval-Augmented Generation (RAG)-based memory to enhance planning and interactive environmental learning (gao2023retrieval, ), but most are designed for simulated environments.
-A key limitation for real-world deployment is the absence of spatial memory, which is critical for building spatial awareness and providing context for planning.
-Additionally, existing multi-module memory systems often incur significant inference latency.
-To overcome these limitations, especially the need for memory that is efficient, spatially grounded, and persistent in dynamic environments, we return to the essence of intelligence — how does the human brain plan, remember, and learn in dynamic environments?
-Inspired by cognitive neuroscience, we have designed RoboMemory, a parallel multi-memory architecture that simulates key functional regions of the brain.
-RoboMemory features a hierarchical and parallelized architecture enabling long-term planning and continuous adaptation.
-Drawing inspiration from cognitive neuroscience (milner1998cognitive, ), RoboMemory comprises four core components (Figure 1): (1) Information Preprocessor (thalamus-inspired) for multimodal sensory integration.
-(2) Comprehensive Embodied Memory System (hippocampus-inspired), which organizes experiential and spatial knowledge through a three-tier structure (long-term, short-term, and sensory memory).
-Within this tiered system, four memory modules: Spatial, Temporal, Episodic, and Semantic operate under a unified, parallel-update paradigm to enable coherent knowledge integration while minimizing latency.
-(3) Closed-Loop Planning Module (prefrontal cortex-inspired) for high-level action sequencing.
-These three modules provide a high-level planner with comprehensive sensory and memorization ability.
-(4) Low-level Executor (cerebellum-inspired), consisting of a VLA-based operation model and a SLAM-based navigation model.
-The Low-level Executor directly controls the robot with low-level control signals to navigate and operate in the real-world environment.
-To verify whether RoboMemory truly addresses the problems of long-horizon planning and interactive learning, we evaluate RoboMemory on EmbodiedBench, a long-horizon planning benchmark (yang2025embodiedbench, ).
-Using Qwen2.5-VL-72B as the base model, RoboMemory improves average success rates by 25% over its base model and 3% over the closed-source state-of-the-art model, Gemini-1.5-Pro.
-In real-world trials, RoboMemory executed diverse tasks twice consecutively: once for environmental familiarization (learning phase) and once for memory-augmented execution (testing phase) without resetting memory.
-The observed performance improvement validates RoboMemory’s capacity for interactive environmental learning.
-We further conduct ablation studies and error analysis to quantify component contributions and identify remaining limitations.
-We summarize our contribution as follows:
-  • We propose a brain-inspired unified embodied memory system, integrating four concurrently updated modules (Spatial, Temporal, Episodic, Semantic) into a single framework.
+In summary, the current VLM–VLA paradigm still lacks (1) a robust spatial and long-term memory mechanism to maintain a consistent view of the environment, and (2) an effective way to acquire and leverage embodied planning experience for generalization.
+To address these challenges, we propose RoboMemory, a brain-inspired memory framework designed to bridge the gap between biological cognitive mechanisms and robotic embodied intelligence.
+Drawing on foundational cognitive theories, we recognize that human memory is inherently organized into a tiered architecture rather than a flat buffer: sensory memory captures transient perceptual inputs [37], short-term working memory handles immediate information manipulation [4, 5], and long-term memory provides persistent storage for episodic events and semantic facts [3, 43].
+Neuroscientific research further reveals that specialized brain regions underpin these functions: the thalamus integrates multimodal sensory inputs, the hippocampus consolidates spatial and episodic experiences into long-term storage [29], the prefrontal cortex regulates high-level planning based on retrieval, and the cerebellum coordinates low-level motor execution.
+Mimicking this biological hierarchy to ensure layered information processing, our recipe for this generalist memory system consists of four key stages (Figure 1): 1) Information Preprocessing (Thalamus-inspired), which integrates multimodal sensory inputs for downstream processing; 2) Comprehensive Embodied Memory (Hippocampus-inspired), which organizes experiential and spatial knowledge through a three-tier structure (sensory, short-term, and long-term).
+Crucially, we design four parallel-update modules (Spatial, Temporal, Episodic, and Semantic) to minimize latency while ensuring data consistency; 3) Closed-Loop Planning (Prefrontal Cortex-inspired), where the planner utilizes retrieved memory to generate high-level action sequences; 4) Low-level Execution (Cerebellum-inspired), which translates plans into precise robot actions using VLA models and SLAM-based navigation.
+While recent frameworks [39, 18, 45, 1, 16, 54, 10] have integrated Retrieval-Augmented Generation (RAG) to enhance planning, they suffer from a critical limitation: most are tailored for simulated environments or rely on flat, text-based retrieval that fails to capture the spatial complexity of the physical world.
+These approaches often lack a hierarchical structure, leading to inefficiencies in long-term knowledge accumulation.
+Moreover, existing multi-module systems [38, 1] typically incur significant inference latency due to sequential memory updates.
+In contrast, our framework leverages a parallel-update paradigm and a hierarchical organization, enabling the agent to maintain a dynamic, spatially-aware worldview in real-time, which is essential for robust interaction in dynamic environments.
+We evaluate RoboMemory on EmbodiedBench, a challenging long-horizon planning benchmark [47].
+Using Qwen2.5-VL-72B as the base model, RoboMemory achieves state-of-the-art results, improving average success rates by 26.5% over the baseline and outperforming the closed-source Claude3.5-Sonnet [2].
+In real-world deployments, our system demonstrates the capacity for continuous improvement: by executing diverse tasks sequentially without memory resets, the agent effectively learns from environmental familiarization.
+Our main contributions are three-fold:
+  • We propose RoboMemory, a brain-inspired, unified embodied memory system that integrates four parallel update modules into a single framework.
 It enables efficient, comprehensive memory operations and coherent knowledge integration, which are critical for interactive environmental learning in real-world embodied scenarios.
-  • We design a retrieval-based incremental update algorithm for real-time evolution of Spatial Knowledge Graphs (KGs).
-By retrieving relevant subgraphs, detecting local inconsistencies, and merging new observations, it ensures efficient, consistent KG maintenance and addresses the scalability bottleneck of previous KG-based methods in embodied settings.
-  • RoboMemory supports interactive environmental learning for real-world physical robots: it enables sequential diverse tasks without memory reset, with experience accumulation driving steady performance improvements, demonstrating practical long-term autonomous learning in physical scenarios.
-2 RoboMemory
+  • We design a retrieval-based incremental update algorithm for the real-time evolution of Spatial Knowledge Graphs (KGs).
+This algorithm bridges the gap between static knowledge representation and dynamic embodied interaction, establishing KGs as a practical and scalable modality for Spatial Memory.
+By retrieving relevant subgraphs, detecting local inconsistencies, and merging new observations, our method ensures efficient maintenance and overcomes the scalability bottlenecks that previously hindered KG-based approaches.
+  • RoboMemory supports interactive environmental learning in real-world environments.
+It enables sequential diverse tasks without memory reset, with experience accumulation driving steady performance improvements, demonstrating practical long-term autonomous learning in physical scenarios.
+Figure 2: (a) Left: The loop where the Planner, Critic, and Comprehensive Embodied Memory interact to adjust plans based on real-time visual inputs.
+Colored text denotes the execution status of actions (success/rejected/replanned).
+(b) Right: Spatial memory maintains a relevance/similarity-updated KG, and Semantic/Episodic memory manages a Vector DB with analogous logic.
+Besides, Temporal memory is implemented as a linear FIFO buffer that stores step-wise summaries generated by the Step Summarizer.
+II Related Work
+II-A VLM/LLM-based Agentic Frameworks in Embodied Tasks
+The rapid advancement of VLMs/LLMs has led to diverse agent frameworks in embodied environments [49, 36, 26].
+Embodied tasks involve partial observability and long-horizon planning, requiring memory systems to retain context.
+Some use time-ordered context buffers for short-term memory (due to VLMs/LLMs’ limited long-context processing) [49, 32]; others adopt experience buffers as long-term semantic memory [16, 35].
+For long-duration tasks, skill libraries serve as procedural memory, with agents accumulating skills via interaction [45, 39].
+However, in real-world settings, the low-level executor may fail to complete the task, making it challenging to construct a reusable, code-based skill library.
+So, explicit procedural memory still needs to be improved in real-world settings.
+Moreover, Recent efforts integrate diverse memories [51, 39, 1] but focus on virtual/GUI environments, leaving real-world multi-modal memory support for long-term planning under-explored.
+II-B Spatial Memory
+Spatial information is critical for embodied agents.
+Prior work has explored three main representations, each with inherent limitations that motivate our high-level abstractions: 2D Semantic Maps encode environments as pixel grids [51].
+As the semantic maps are flattened, they discard 3D relationships (e.g., “above/beneath”) and precise geometry, limiting expressiveness for manipulation tasks requiring fine-grained spatial reasoning.
+3D Point Clouds preserve accurate 3D information [52].
+However, a 3D point cloud only encodes low-level spatial information.
+We need to put in extra effort to help agents understand 3D Point Clouds.
+Semantic Scene Graphs [38, 28, 19, 48, 20] offering a compact, high-level representation directly amenable to planning.
+However, existing methods suffer from two key drawbacks: (1) Most previous methods adopt rigid tree-like structures, which are easy to update with new information, but cannot model non-hierarchical or cyclic spatial configurations.
+(2) Some previous methods try to use general graphs to store various spatial information.
+[20] However, their design often lacks efficient dynamic update mechanisms—adding or modifying relations typically requires expensive full-graph recomputation, creating a scalability bottleneck for real-time embodied interaction.
+For clarity, we summarize the differences among different memory systems in the previous work.
+The comparison is shown in Table III in the appendix.
+Additionally, we provide more ablation studies in Appendix -A.
+III RoboMemory
 RoboMemory is a hierarchical embodied agent system that equips robots with three core memory capabilities: historical interaction logs, dynamically updated spatial layouts, and accumulated task knowledge.
-As illustrated in Figure 2, each iteration, RoboMemory follows a process of “Perception – Memory – Retrieval – Planning – Execution” process, ensuring that the agent continuously calibrates its memory and behavior in dynamic environments.
+As illustrated in Figure 2, in each iteration, RoboMemory follows a process of “Perception – Memory – Retrieval – Planning – Execution” process, ensuring that the agent continuously calibrates its memory and behavior in dynamic environments.
 First, the information preprocessor converts multimodal sensor inputs into a textual summary of the current scene, which serves as the primary input to the Comprehensive Embodied Memory.
 Next, the Comprehensive Embodied Memory updates its internal representations, including action histories, object locations, and experiential knowledge.
-After information is updated, the memory system retrieves contextually relevant entries to inform the Closed-Loop Planning Module.
+After updating the information, the memory system retrieves contextually relevant entries to inform the Closed-Loop Planning Module.
 Then, leveraging this contextual memory, the Closed-Loop Planning Module generates high-level, text-based action instructions.
-Finally, these commands are dispatched to low-level executors, who will directly control the robot and complete the instructions.
-Figure 2: RoboMemory architecture.
-(a) Left: Parallel Step Summarizer and Query Generator generate updates/queries for Comprehensive Embodied Memory.
-These memories enable Closed-Loop Planning for tasks like “slice and pick up the apple”—the Planner generates plans, while the Critic and memories adjust decisions via feedback from visual inputs/results.
-(b) Right: Spatial memory maintains a relevance/similarity-updated KG, and Semantic/Episodic memory manages a Vector DB with analogous logic.
-Besides, Temporal memory is implemented as a linear FIFO buffer that stores step-wise summaries generated by the Step Summarizer.
-2.1 Information Preprocessor
-At each time step $i$, RoboMemory receives a visual observation $\mathcal{O}_{i}$: an RGB frame (in simulation) or a short video clip (on physical robots), representing the agent’s observations.
-Since raw visual data is unsuitable for direct use in memory construct and retrieval, RoboMemory first employs an information preprocessor to convert multimodal observations into textual representations, thereby providing a semantic interface for subsequent memory and planning modules.
-The information preprocessor executes two Vision-Language Models (VLMs) in parallel: (1) Step summarizer $\mathcal{S}$: Transforms $\mathcal{O}_{i}$ into a concise textual description $s_{i}$ of the just-executed action.
-The string $s_{i}$ is stored in the system’s working memory.
-(2) Query generator $\mathcal{Q}$: Derives a query $q_{i}$ from the same observation $\mathcal{O}_{i}$, which is used to probe long-term memory for relevant episodes.
+Finally, these commands are dispatched to low-level executors, which directly control the robot and complete the instructions.
+The execution process is demonstrated in Algorithm 1 in the appendix.
+III-A Information Preprocessor
+At each time step $t$, RoboMemory receives a visual observation $\mathcal{O}_{t}$: an RGB frame (in simulation) or a short video clip (on physical robots), representing the agent’s observations.
+Since raw visual data is unsuitable for direct use in memory construction and retrieval, RoboMemory first employs an information preprocessor to convert multimodal observations into textual representations, thereby providing a semantic interface for subsequent memory and planning modules.
+The information preprocessor executes two Vision-Language Models (VLMs) in parallel: (1) Step summarizer $\mathcal{S}$: It transforms $\mathcal{O}_{t}$ into a concise textual description $s_{t}$ of the just-executed action.
+The string $s_{t}$ is stored in the system’s working memory.
+(2) Query generator $\mathcal{Q}$: It derives a list of queries $q_{t}=[q_{t}^{(1)},q_{t}^{(2)},\dots,q_{t}^{(N)}]$ from the same observation $\mathcal{O}_{t}$.
+Each query $q^{(i)}_{t}$ is a natural language-based query.
+These queries are used to query information from the memory system that may be useful.
 Together, $\mathcal{S}$ and $\mathcal{Q}$ provide a swift, text-based interface between raw sensory data and provide basic information in each iteration for RoboMemory’s Comprehensive Embodied Memory System.
-2.2 Comprehensive Embodied Memory System
-We present a high-level memory system designed to enhance embodied agents’ long-horizon planning and enable interactive learning in dynamic, partially observable environments.
-Built on RAG and KG technologies (gutierrez2024hipporag, ), the system comprises Spatial-Temporal Memory and Long-Term Memory.
-Spatial-Temporal Memory continuously encodes object relationships and environmental layouts, as well as detailed task execution traces.
-These records inform Long-Term Memory, which goes through spatial-temporal information collected in each task, capturing what strategies succeeded or failed and relevant environmental facts (Figure 2).
-The Comprehensive Embodied Memory module receives summarized sensory information from the information preprocessor to update both memory components in real time.
-Critically, all updates and retrievals are performed in parallel, ensuring efficient operation even under multiple active memory modules.
-2.2.1 Spatial-Temporal Memory System
-Due to the dynamic and partially observable nature of real embodied environments, completing long-term tasks requires dynamically maintaining a spatial-temporal memory system.
-The system records the agent’s interaction history with the environment and maintains a continuously updated spatial model of the current surroundings.
-To achieve these two purposes, we design spatial memory and temporal memory.
-The combination of spatial and temporal memory provides the embodied agent with comprehensive contextual awareness, supporting effective planning and interactive environmental learning in real-world settings.
+III-B Comprehensive Embodied Memory
+To address the long-term memory limitations in current embodied agent frameworks, we propose the Comprehensive Embodied Memory System.
+This system consists of multiple memory modules.
+We denote the memory system containing $L$ distinct modules as $M_{t}=[M_{t}^{(1)},M_{t}^{(2)},...,M_{t}^{(L)}]$, where $M_{t}$ represents the memory stored at step $t$, and $M_{t}^{(l)}$ denotes the $l$-th memory module.
+Generally, the update and retrieval process at iteration $t$ of a general memory is shown below:
+$M_{t}=\mathcal{U}(M_{t-1},s_{t})$ (1)
+$r_{t}=\mathcal{R}(M_{t},q_{t})$ (2)
+First, we update memory modules with memory update algorithm $\mathcal{U}$, where we update the previous memory $M_{t-1}$ using the latest summarization $s_{t}$.
+Then, with updated $M_{t}$, we use memory retrieve algorithm $\mathcal{R}$ to retrieve the information that is useful for the planning module.
+In $\mathcal{R}$, we use queries $q_{t}$ to query $M_{t}$, yielding retrieval results from each module: $r_{t}=[r_{t}^{(1)},r_{t}^{(2)},...,r_{t}^{(L)}]$.
+These results are then passed to the planning module, which helps it plan future movements.
+However, sequentially updating and retrieving from $L$ modules would be a slow process.
+Therefore, we parallelize these steps across all modules, which significantly enhance the efficiency of the memory system.
+The memory system consists of four distinct modules ($L=4$): Temporal Memory, Spatial Memory, Semantic Memory, and Episodic Memory.
+For efficiency, all memory modules are updated and retrieved in parallel.
+Thus, even with multiple modules, the system remains highly efficient.
+Functionally, inspired by cognitive psychology [27], our modules handle memory at different levels.
+In cognitive psychology, memory is divided into Sensory Memory, Short-term Memory, and Long-term Memory.
+Mirroring this hierarchy, our modules are organized as follows.
+First of all, the step summarizer $\mathcal{S}$ summarizes the agent’s interactions with the environment at each iteration.
+It acts as Sensory Memory.
+Secondly, Temporal Memory and Spatial Memory function as short-term memory.
+These two memories will be updated at each iteration.
+They are designed to store the information of sensory memory in every iteration.
+For Temporal Memory, we record the agent’s action history sequentially, while for Spatial Memory, we dynamically record the spatial relationships between different objects in the environment based on Sensory Memory.
+These memories can provide a relatively long and detailed history of the current task for the Closed-Loop Planner to make a future plan.
+Thirdly, Semantic Memory and Episodic Memory serve as Long-term Memory.
+They are updated only when meaningful information arises (e.g., after task completion).
+These memories store highly abstract knowledge, not limited to the current task, but synthesized from past experiences.
+This knowledge—factual, event-based, and experiential—improves the agent’s future task performance.
+It is the source of RoboMemory’s interactive learning capability.
+We now detail each module.
 Temporal Memory.
-Interactions between the robot and the environment must be recorded.
-For such temporally sequential memory, a simple structure suffices: a sequential buffer with automatic summarization triggered upon reaching capacity.
-The buffer tracks the high-level actions from the Embodied Agent, how the low-level executor actually controls the robot, and the environmental feedback received.
+In the Temporal Memory, we record interactions between the robot and the environment (i.e., Sensory Memory) of each iteration sequentially.
+This information can provide the embodied agent with simple awareness of “What I have done”.
+For such temporally sequential memory, a simple structure is sufficient: a sequential buffer with automatic summarization triggered when the record sequence reaches its capacity.
 In a specific design, temporal memory can store up to $N$ interaction summaries, each generated by an information preprocessor.
-When the buffer is full, rather than discarding old entries outright, we compress the oldest $N$ steps into a single summarized entry using a VLM, which is then reinserted at the front of the buffer, ensuring continuous context retention without unbounded growth.
+When the buffer is full, we compress the oldest $N$ steps into a single summarized entry using a VLM, which is then reinserted at the front of the buffer, ensuring continuous context retention without unbounded growth.
+However, the information from previous memories will gradually be lost as we summarize it multiple times.
+For retrieval, we provide all existing memory in text to downstream modules.
 Spatial Memory.
-Current spatial memory approaches often rely on RGB-D cameras to reconstruct 3D point clouds (zhang2023building, ; chang2023goat, ).
-However, such representations are overly detailed for high-level planning in embodied agents — precise geometric relationships (e.g., exact distances between objects) are unnecessary.
-Moreover, point clouds are difficult to update efficiently in dynamic scenes.
-To address these problems, we advocate for an abstracted spatial memory that captures object-level entities and their spatial relationships.
-Inspired by gutierrez2024hipporag , we represent this memory as a KG: objects become vertices, and spatial relations are encoded as edges.
+The spatial memory is designed to dynamically record the high-level spatial relationships of different entities in the environment.
+However, current spatial memory approaches often rely on RGB-D cameras to reconstruct 3D point clouds [51, 9].
+These representations are too detailed for high-level planning in embodied agents.
+For example, precise geometric relationships (e.g., exact distances between objects) are unnecessary.
+To address these problems, we use a dynamic KG to store high-level spatial information: objects and positions in the environment become vertices of the KG, and spatial relations between objects or positions are encoded as edges.
 The KG focuses on high-level spatial relations (e.g., “cup on table”, “key left of drawer”).
-Unlike geometric representations that encode precise coordinates, which are often irrelevant for high-level planning, our KG focuses on semantically meaningful, task-relevant relations.
-This abstraction enhances the agent’s spatial reasoning capability in dynamic environments.
-In our memory system, spatial memory is a dynamic KG storing high-level spatial knowledge.
-To overcome the limited spatial reasoning of embodied agents in dynamic environments, we use a KG.
-We introduce a retrieval-driven, incremental KG update algorithm that maintains a locally modifiable, globally consistent, and dynamically adaptive spatial memory.
+By these settings, spatial KG focuses on semantically meaningful, task-relevant relations.
+This spatial information enhances the agent’s spatial reasoning capability in dynamic environments.
+However, as related work shows, most KG construction algorithms are designed for static long content.
+This does not meet the demand of using KG as spatial memory for the agent.
+The KG needs to update efficiently in response to new information.
+To address this issue, we introduce a retrieval-driven, incremental KG update algorithm that maintains a locally modifiable, globally consistent, and dynamically adaptive spatial memory.
 As illustrated in the right panel of Figure 2, the update process proceeds in four steps: (1) retrieves the most relevant sub-KG around new observations.
@@
 (3) Detects and resolves conflicts between newly extracted relations and existing ones (e.g., “cup on table” vs.
-‘cup in drawer”) using a VLM-based resolver, which decides whether to add, delete, or modify edges.
+“cup in drawer”) using a VLM-based resolver, which decides whether to add, delete, or modify edges.
 (4) Merges back and prunes isolated vertices.
-Moreover, our retrieval-based incremental update algorithm comes with provable efficiency guarantees: for a KG with $n$ vertices and maximum degree $D$, the number of vertices processed per update is bounded by $O(D^{K})$, where $K$ is the retrieval hop distance (see Appendix C.3 for formal analysis).
-Further architectural and implementation details are provided in Appendix C.1.
-2.2.2 Interactive Environmental Learning System
-Although episodic memory can record “what happened,” it cannot directly answer “why it succeeded/failed” or “what should be done next time.” Drawing on cognitive psychology’s classification of human long-term memory, we divide our system into episodic and semantic memory: the former records agent-environment interaction histories, while the latter extracts experiential insights to support long-term task reasoning.
-This update process mirrors human daily experience consolidation during sleep (maboudi2024retuning, ).
-Episodic Memory: It captures task-level interactions, accounting for temporal interdependencies between sequential tasks in the same environment.
-The agent must memorize what it has done before to complete future tasks.
-Moreover, task-level interactions can be a reference that may help the agent improve its plan in the future.
-Semantic Memory: It accumulates step-by-step action usage experiences (based on invoked actions and outcomes) to inform action arrangement.
-Post-task, it summarizes temporal memory, distilling successes from completed tasks and identifying failure causes/improvement strategies from unsuccessful ones, thus enabling both action-level and task-level learning across both task and action levels.
-In implementation, both episodic and semantic memory share the same RAG framework consisting of an extractor, updater, and RAG storage (each entry is a memory entity).
-Post-task, the Extractor summarizes the task’s Spatial-Temporal Memory into a new memory entity.
-The RAG then retrieves similar existing entities (old information) from the RAG.
-Then, the Updater deletes, adds, or updates old memory entities according to new information.
-After that, we write the updated memory entities back to the RAG.
-Because we only update memory entities similar to new information, efficiency is ensured by restricting updates to old memory entities relevant to the new entity instead of all memory entities stored in the RAG.
-2.3 Closed-Loop Planning Module for Dynamic Environment
+Moreover, our retrieval-based incremental update algorithm is accompanied by provable efficiency guarantees.
+For a KG with $n$ vertices and maximum degree $D$, the number of vertices processed per update is bounded by $O(D^{K})$, where $K$ is the retrieval hop distance (see Appendix -D for analysis).
+Further implementation details are provided in Appendix -C1.
+Semantic Memory.
+In cognitive psychology, semantic memory stores time-independent facts.
+These facts are stable, update slowly, and require long-term retention.
+In RoboMemory, semantic memory records task-relevant experiences and environmental knowledge during execution.
+This information can help RoboMemory adapt to new environments or tasks.
+This information is highly abstract and does not need to be updated frequently.
+In RoboMemory, semantic memory updates when new information is encountered during execution, for example, after completing a subtask or encountering important information.
+To store and update memory efficiently, we design a memory management system based on a vector database (DB).
+In the vector DB, each experience/fact is described in natural language (denoted as a memory item).
+Each memory item is converted into a semantic vector for querying.
+For dynamic updates, we adapt a basic method from prior work [12].
+However, we adopt it into the embodied environment.
+As shown in the bottom-right of Figure 2, the semantic memory update algorithm involves two VLM-based modules.
+Firstly, the Long-term Memory Creator generates new memory items based on short-term memory.
+We retrieve the top-$S$ most similar existing memory items from existing memory via cosine similarity.
+A VLM-based updater then compares new and existing items to decide whether to: add the new item, update an existing item, remove an outdated item, or perform Noop (if redundant).
+Since updates only involve a maximum of $S$ previous memory items and the update process is parallelized across all memory modules, this update method ensures that semantic memory remains efficient even as the database grows.
+For retrieval, we use the same process as a traditional vector DB.
+We use queries to extract top-$N$ relevant information for downstream modules.
+Episodic Memory.
+In cognitive psychology, Episodic Memory is another important part of long-term memory.
+It can store task-specific execution summaries (i.e., “autobiographical” records of past attempts).
+In RoboMemory, the Episodic Memory module is responsible for recording every interaction trajectory it has gone through, including the sequence of actions the robot did and the feedback from the environment.
+The trajectory information can help to improve the planning ability of RoboMemory.
+For example, if a trajectory for completing a similar task is stored in episodic memory, it can guide the agent in completing the current task.
+As the agent only needs to follow the successful trajectory in the memory, it can reduce hallucinations or errors in the VLM planner.
+Technically, Episodic Memory shares the same storage and VLM-driven vector DB update mechanism as Semantic Memory, ensuring consistent architectural design.
+III-C Closed-Loop Planning Module
 The Closed-Loop Planning Module integrates information about the current task provided by the Spatial-Temporal Memory, Semantic and Episodic information recorded in long-term memory, and current observations to perform action planning.
 Each action is planned and passed on to the low-level executor for execution.
-To enable closed-loop control in embodied environments, the Closed-Loop Planning Module adopts the Planner-Critic mechanism (lei2025clea, ), which consists of the planner and the critic module.
+To enable closed-loop control in embodied environments, the Closed-Loop Planning Module adopts the Planner-Critic mechanism [25], which consists of the planner and the critic module which is powerd by same VLM model with different prompts.
+We denote the planner module as $\mathcal{P}$, while the critic module as $\mathcal{C}$.
 For each planning step, the planner generates a long-term plan consisting of multiple steps.
@@
 In the original mechanism, the first step of the action sequence output by the Planner is evaluated by the Critic before execution, which can lead to an infinite loop: if the Critic always demands replanning, no action will ever be executed.
-To address this, we modified the Planner-Critic mechanism so that the first step is not evaluated by the Critic.
+To address this, we modify the Planner-Critic mechanism so that the first step is not evaluated by the Critic.
 This ensures that even if the Critic persistently demands replanning, the RoboMemory will still execute actions.
-2.4 Low-level Executor
-The RoboMemory framework is a two-layer hierarchical agent framework.
-This design enables RoboMemory to accomplish longer-term tasks in the real world.
+The detailed algorithm is shown in Algorithm 1 in the appendix.
+III-D Low-Level Executor
+The RoboMemory is a two-layer hierarchical agent framework that accomplishes longer-term tasks in the real world.
 The upper layer is responsible only for high-level planning, while the Low-level Executor carries out the actions planned by the upper layer in the real environment.
-We employ a LoRA-finetuned VLA model, $\pi_{0}$ (hu2022lora, ; black2024pi_0, ), to generate manipulation actions, and a SLAM-based navigation model for locomotion.
+We employ a LoRA-finetuned VLA model, $\pi_{0}$ [21, 8], to generate manipulation actions, and a SLAM-based navigation model for locomotion.
 The low-level executor then translates high-level actions planned by RoboMemory into concrete arm and chassis movements in the real world.
-Table 1: Comparison of Success Rates (SR) and Goal Condition Success Rates (GC) across difficulty levels between RoboMemory and baseline methods on EB-ALFRED.
-Method Type Avg. Base Long
-SR GC SR GC SR GC
+TABLE I: Comparison of Success Rates (SR) and Goal Condition Success Rates (GC) across difficulty levels (Base/Long) on EB-ALFRED and EB-Habitat benchmarks. Values are reported in percentages (%).
+Method Type EB-ALFRED EB-Habitat
+Average Base Long Base Long
+SR GC SR GC SR GC SR GC SR GC
 Single VLM-Agents
-GPT-4o Closed-source 59.0 % 68.3 % 64.0 % 74.0 % 54.0 % 62.5 %
-GPT-4o-mini 17.0 % 32.4 % 34.0 % 47.8 % 0.0 % 17.0 %
-Claude-3.5-Sonnet 62.0 % 63.3 % 72.0 % 72.0 % 52.0 % 54.5 %
-Gemini-1.5-Pro 64.0 % 69.7 % 70.0 % 74.3 % 58.0 % 65.0 %
-Gemini-2.0-flash 60.0 % 63.9 % 62.0 % 65.7 % 58.0 % 62.0 %
-Llama-3.2-90B-Vision-Ins Open-source 27.0 % 33.9 % 38.0 % 43.7 % 16.0 % 24.0 %
-InternVL2.5-78B 40.0 % 45.7 % 38.0 % 42.3 % 42.0 % 49.0 %
-InternVL3-78B 37.0% - 38.0% - 36.0% -
-InternVL2.5-38B 31.0% 36.9% 36.0% 37.3% 26.0% 36.5%
-Qwen2.5-VL-72B-Ins 42.0 % - 50.0 % - 34.0 % -
+GPT-4o Closed-source 67.0 74.9 64.0 74.0 54.0 62.5 86.0 90.7 64.0 72.2
+GPT-4o-mini 30.5 40.9 34.0 47.8 0.0 17.0 74.0 77.5 14.0 21.3
+Claude-3.7-Sonnet 68.5 - 68.0 - 70.0 - 90.0 - 46.0 -
+Claude-3.5-Sonnet 69.5 71.8 72.0 72.0 52.0 54.5 96.0 97.5 58.0 63.3
+Gemini-1.5-Pro 68.0 73.3 70.0 74.3 58.0 65.0 92.0 92.5 52.0 61.2
+Gemini-2.0-flash 57.0 61.5 62.0 65.7 58.0 62.0 82.0 82.0 26.0 36.2
+Llama-3.2-90B-Vision-Ins Open-source 40.5 46.6 38.0 43.7 16.0 24.0 94.0 94.5 14.0 24.3
+InternVL2.5-78B 47.0 52.9 38.0 42.3 42.0 49.0 80.0 82.0 28.0 38.2
+InternVL2.5-38B 37.5 42.6 36.0 37.3 26.0 36.5 60.0 61.5 28.0 35.0
+InternVL3-78B 49.5 - 38.0 - 36.0 - 84.0 - 40.0 -
+Qwen2.5-VL-72B-Ins 44.0 - 50.0 - 34.0 - 74.0 - 18.0 -
 VLM-Agent Frameworks
-Voyager (Qwen2.5-VL-72B-Ins) Baselines 44.0% 63.7% 56.0% 73.2% 32.0% 54.2%
-Reflexion (Qwen2.5-VL-72B-Ins) 29.0% 43.2% 48.0% 54.0% 10.0% 33.0%
-Cradle (Qwen2.5-VL-72B-Ins) 43.0% 54.6% 54.0% 67.9% 32.0% 41.0%
-RoboMemory (Qwen2.5-VL-72B-Ins) Ours 67.0 % 78.4 % 68.0 % 75.5 % 66.0 % 81.3 %
-3 Experiments
-3.1 Benchmarks
-To evaluate the task planning ability of RoboMemory, we select a subset of the EB-ALFRED benchmark from EmbodiedBench(yang2025embodiedbench, ).
-We selected the Base and Long subsets because they aim to test the agent’s planning ability.
-The Base and Long subset comprises 100 tasks for complex embodied tasks.
-The EB-ALFRED environment provides a visually-grounded operational setting that closely mimics real-world conditions (see Appendix D for environment details), enabling direct comparison with established baselines.
-Additionally, we have also evaluated RoboMemory’s capabilities on the EB-Habitat benchmark (yang2025embodiedbench, ). For detailed results, please refer to Appendix B.
+Voyager (Qwen2.5-VL-72B-Ins) Baselines 46.5 66.4 56.0 73.2 32.0 54.2 76.0 87.0 22.0 51.0
+Reflexion (Qwen2.5-VL-72B-Ins) 38.3 51.1 48.0 54.0 10.0 33.0 80.0 84.2 15.0 33.0
+Cradle (Qwen2.5-VL-72B-Ins) 44.5 57.0 54.0 67.9 32.0 41.0 62.0 67.0 30.0 52.1
+RoboOS (Qwen2.5-VL-72B-Ins) 25.5 33.0 32.0 38.4 12.0 17.6 38.0 47.8 20.0 28.2
+RoboOS (RoboBrain2-32B) 20.0 25.7 32.0 37.2 8.0 13.2 28.0 34.8 12.0 17.4
+RoboMemory (Qwen2.5-VL-72B-Ins) Ours 70.5 79.7 68.0 75.5 66.0 81.3 86.0 88.0 62.0 74.0
+IV Experiments
+IV-A Benchmarks
+To evaluate RoboMemory’s task planning ability, we select a subset of the EmbodiedBench EB-ALFRED and EB-Habitat benchmark [47].
+We select the Base and Long subsets because they aim to test the agent’s planning ability.
+The Base and Long subsets of the two benchmarks comprise 200 tasks for complex embodied tasks.
+The EB-ALFRED and EB-Habitat benchmarks provide a visually grounded operational setting that closely mimics real-world conditions (see Appendix -E for environment details), enabling direct comparison with established baselines.
+We set temperature = 0 to avoid randomness during the experiment, aligning with the EmbodiedBench setting.
+Each task is executed once.
 Moreover, we set up an environment to test the interactive environmental learning ability of RoboMemory in the real world.
-3.2 Settings & Baselines
+IV-B Baselines and Metrics
 To facilitate comparisons, we consider two types of baselines.
@@
 We compare their performance with RoboMemory.
-For closed source VLMs, we choose GPT-4o and GPT-4o-mini (GPT-4o-mini, ; hurst2024gpt, ), Claude3.5-Sonnet (Claude-3.5-Sonnet, ), Gemini-1.5-Pro and Gemini-2.0-flash (team2024gemini, ; Gemini2.0, ).
-For open source VLMs, we choose Llama-3.2-90B-Vision-Ins (llama3.2, ), InternVL-2.5-78B/28B (chen2024expanding, ), InternVL-3-72B (zhu2025internvl3, ), and Qwen2.5-VL-72B-Ins (bai2025qwen2, ).
-Secondly, we choose three agent frameworks: (1) Reflexion (shinn2024reflexion, ), which introduces a simple long-term memory and a self-reflection module.
+For closed source VLMs, we choose GPT-4o and GPT-4o-mini [31, 23], Claude3.5-Sonnet and Claude-3.7-Sonnet [2], Gemini-1.5-Pro and Gemini-2.0-flash [41, 15].
+For open source VLMs, we choose Llama-3.2-90B-Vision-Ins [30], InternVL-2.5-78B/28B [11], InternVL-3-72B [55], and Qwen2.5-VL-72B-Ins [6].
+Secondly, we choose three agent frameworks: (1) Reflexion [35], which introduces a simple long-term memory and a self-reflection module.
 Reflexion uses the self-reflection module to summarize experiences as long-term memory, thereby enhancing the model’s capabilities.
-(2) Voyager (wang2023voyager, ), which utilizes a skill library as its procedural memory, is a widely used baseline for embodied agent planning.
-(3) Cradle (tan2024cradle, ), which proposes a general agent framework with episodic and procedural memory and gains good performances at various multi-model agent tasks.
-In our experiments, each agent framework is tested using Qwen2.5-VL-72b-Ins (qwen2.5, ).
+(2) Voyager [45], which utilizes a skill library as its procedural memory, is a widely used baseline for embodied agent planning.
+(3) Cradle [39], which proposes a general agent framework with episodic and procedural memory and gains good performances at various multi-model agent tasks.
+(4) RoboOS [38], which proposes an embodied agent framework that consists of a Hierarchy Scene-Graph based Spatial Memory.
+In our experiments, each agent framework is tested using Qwen2.5-VL-72b-Ins [42] with temperature set as 0.
+For the RoboOS framework, we test it on RoboBrain2-32B [40], where the RoboBrain2-32B model is designed for the RoboOS framework.
 The Qwen2.5-VL-72b-Ins represents a high-performing open-source alternative.
-Notably, the Qwen2.5-VL-72b-Ins demonstrates performance comparable to advanced closed-source VLMs in several benchmark tasks (white2024livebench, ).
-We use the Qwen3-Embedding model (zhang2025qwen3, ) to create embedding vectors for RAGs in RoboMemory.
+Notably, the Qwen2.5-VL-72b-Ins demonstrates performance comparable to advanced closed-source VLMs in several benchmark tasks [46].
+We use the Qwen3-Embedding model [53] to create embedding vectors for RAGs in RoboMemory.
 For the Low-level Executor, since EB-ALFRED provides high-level action APIs, we use the Low-level executor provided by EmbodiedBench instead of the VLA-based method.
@@
 $GC=\mathbb{E}_{x\in\mathcal{X}}\left[\frac{SCN_{x}}{GCN_{x}}\right]$ (4)
-Where $\mathcal{X}$ denotes the test subset, and $x$ represents a test task.
+In above formulas, $\mathcal{X}$ denotes the test subset, and $x$ represents a test task.
 The success condition number ($SCN_{x}$) refers to the number of conditions the agent has accomplished, while the global condition number ($GCN_{x}$) indicates the total number of conditions required for task completion.
 The task is considered successful if $SCN_{x}=GCN_{x}$.
-Figure 3: Efficiency improvement of Comprehensive Embodied Memory System
-Table 2: Ablation Study on RoboMemory’s Success Rate (SR)
-Method Avg. Base Long
-RoboMemory 67% 68% 66%
-- w/o critic 55 % 60 % 50%
-- w/o spatial memory 47 % 52 % 42 %
-- w/o long-term memory 57% 66% 48%
-- w/o episodic memory 62% 68% 56%
-- w/o semantic memory 58% 66% 50%
-3.3 Main results
-As shown in Table 1, our model achieves significant improvements over both single VLM agents and Agent frameworks on the EB-ALFRED.
-Compared to the SOTA Single VLM-Agent model, Gemini1.5-Pro, RoboMemory with Qwen2.5-VL-72B-Ins backbone improves the average SR by 3% and GC by 8.7%.
+IV-C Main Results
+Figure 3: Efficiency improvement of Comprehensive Embodied Memory System.
+As shown in Table I, our model achieves significant improvements over both single VLM agents and Agent frameworks on the EB-ALFRED and EB-Habitat.
+Compared to the SOTA Single VLM-Agent model, Claude3.5-Sonnet, RoboMemory with Qwen2.5-VL-72B-Ins backbone improves the average SR by 1% and GC by 7.9%.
 This demonstrates RoboMemory’s superiority over single VLM-Agents, proving that an Agent framework with open-source models can outperform closed-source SOTA models.
 Furthermore, when tested against other VLM-Agent frameworks, RoboMemory also shows substantial gains.
-This is because, unlike other frameworks, RoboMemory’s brain-like memory system provides embodied models with more accurate and persistent contextual information.
+This is because, unlike other agent frameworks, RoboMemory’s brain-like memory system provides embodied models with more accurate and persistent contextual information.
 Additionally, the Planner-Critic mechanism provides a closed-loop planning ability, which helps the RoboMemory gain better performance in long-term tasks.
@@
 And it is more robust when encountering unexpected situations.
-3.4 Efficiency Analysis
-To evaluate the efficiency of the Comprehensive Embodied Memory module, we tasked RoboMemory with executing 10 long-horizon tasks, each comprising approximately 50 steps.
+IV-D Efficiency Analysis
+To evaluate the efficiency of the Comprehensive Embodied Memory module, we tested RoboMemory with executing 10 long-horizon tasks, each comprising approximately 50 steps.
+All testing is under the same hardware and software settings.
 We exclusively measured the wall-clock time consumed by memory update and retrieval operations.
@@
 This demonstrates the critical efficiency gains afforded by parallelization across the memory architecture.
-3.5 Ablation Studies
+TABLE II: Ablation study of RoboMemory’s Success Rate (SR) on EB-ALFERD subset.
+Method (EB-ALFRED) Avg. Base Long
+RoboMemory 67% 68% 66%
+- w/o critic 55 % 60 % 50%
+- w/o spatial memory 47 % 52 % 42 %
+- w/o episodic memory 62% 68% 56%
+- w/o semantic memory 58% 66% 50%
+- w/o long-term memory 57% 66% 48%
+Figure 4: Visualization of the experimental environment.
+Figure 5: Real-world experiment results. “Qwen” denotes Qwen2.5-VL-72B-Ins; “RoboBrain” denotes RoboBrain2.0-32B.
+IV-E Ablation Studies
 We used the full Base and Long Subset from EB-ALFRED to validate RoboMemory’s effectiveness.
@@
 We use the success rate as our metric.
-Results are shown in Table 2.
-Long-term Memory Adding long-term memory significantly improved RoboMemory’s success rate.
+Results are shown in Table II.
+Long-term Memory: Adding long-term memory significantly improves RoboMemory’s success rate.
 The experiment shows that it enables interactive environmental learning while attempting to complete tasks.
-The semantic memory learns low-level skills’ properties, such as in what circumstances an action may fail.
-The temporal memory records all task attempts (successful/failed), providing valuable experience at the task level and giving insight into how to complete a task successfully.
-This helps the RoboMemory predict action outcomes and avoid ineffective attempts.
+Semantic memory learns the properties of low-level skills, such as in what circumstances an action may fail.
+The episodic memory records all task attempts (successful/failed), providing valuable experience at the task level and giving insight into how to complete a task successfully.
+This helps the RoboMemory predict the outcomes of actions and avoid ineffective attempts.
 This ability indicates that the RoboMemory has an interactive environmental learning capability.
-Spatial Memory Spatial memory is crucial for embodied agents, especially given that current pretrained VLMs have limited spatial understanding ability.
+Additionally, we do ablations for semantic memory and episodic memory solely.
+The result shows that those memory modules significantly improve the capability of the performance of RoboMemory.
+Spatial Memory: Spatial memory is crucial for embodied agents, especially given that current pretrained VLMs have limited spatial understanding ability.
 Our novel dynamic KG update algorithm enables KG-based spatial memory in dynamic environments.
 This spatial reasoning helps RoboMemory handle partially observable embodied settings.
-Critic Module Table 2 shows performance without the critic module (55% vs 67% with full system).
+Critic Module: Table II shows performance without the critic module (55% vs 67% with full system).
 This drop highlights how the critic’s closed-loop planning adapts to dynamic environments.
 It helps RoboMemory recover from failures faster and handle unexpected situations better.
-Figure 4: Visualization of the experimental environment.
-Figure 5: The improvement of RoboMemory after learning in the real world.
-3.6 Real-world Robot Deployment
+IV-F Real-World Robot Deployment
 To evaluate RoboMemory’s interactive environmental learning capability in the real world, we designed a kitchen environment inspired by EB-ALFRED and EB-Habitat.
@@
 Due to search and error recovery, the robot often exceeds 20 steps per task.
-Additional hardware experiment details are in Appendix D.
-To test the interactive environmental learning ability of RoboMemory, we run each task twice without clearing long-term memory between attempts.
-Success rates for first and second attempts are shown in Figure 5.
+Additional hardware experiment details are in Appendix -E.
+To test the interactive environmental learning ability of RoboMemory, we ran each task twice without clearing long-term memory between attempts.
+Meanwhile, we compared RoboMemory against previous SOTA on real-world experiments RoboOS as baselines.
+The success rates for first and second attempts and different settings of RoboOS are shown in Figure 5.
 The second attempt showed significantly higher success rates.
@@
 These analyses guide future decisions.
-Detailed examples demonstrating these capabilities and further discussions are provided in Appendix E.
-Moreover, we observed a significant drop in task success rates when deploying the agent with the Low-level Executor in real-world environments.
-This performance degradation primarily stems from the executor’s inherent limitations: (1) The VLA model exhibits unreliable instruction-following capabilities, frequently failing during grasping actions or selecting incorrect objects; (2) Pre-trained VLM models demonstrate inadequate video understanding - while capable of recognizing static objects, they struggle to interpret dynamic visual information such as action failures or state changes.
+Detailed examples demonstrating these capabilities and further discussions are provided in Appendix -F.
+Moreover, we observe a significant drop in task success rates when deploying the agent with the Low-level Executor in real-world environments.
+This performance degradation primarily stems from the executor’s inherent limitations: (1) The VLA model exhibits unreliable instruction-following capabilities, frequently failing during grasping actions or selecting incorrect objects; (2) Pre-trained VLM models demonstrate inadequate video understanding capability.
+They are struggle to interpret dynamic visual information such as action failures or state changes.
 These limitations collectively contribute to the reduced performance compared to simulated environments.
-4 Conclusion and Future Work
-In summary, RoboMemory, a brain-inspired multi-memory framework, facilitates long-horizon planning and interactive environmental learning in real-world embodied systems by addressing key challenges such as memory latency, task correlation capture, and planning loops.
+V Conclusion
+This paper propose RoboMemory, a brain-inspired multi-memory framework, facilitating long-horizon planning and interactive environmental learning in real-world embodied systems by addressing key challenges such as memory latency, task correlation capture, and planning loops.
 Experiments on EmbodiedBench demonstrate that RoboMemory outperforms state-of-the-art closed-source VLMs and agent frameworks, with ablation studies confirming the critical roles of the Critic module and spatial/long-term memory.
@@
 Despite limitations arising from reasoning errors and executor dependence, RoboMemory provides a foundation for generalizable, memory-augmented agents, with future work aimed at refining reasoning and enhancing execution robustness.
-A notable open challenge in hierarchical embodied agents, including RoboMemory, lies in the interface between high-level planners and low-level executors.
-Existing frameworks typically rely on language instructions to convey actions, yet some execution details (e.g., precise grasp points) are difficult to describe textually and are better captured through other modalities, such as vision.
-While our current work emphasizes long-term planning and interactive learning, future research may improve generalization by developing richer multimodal interactions between the agent and executor.
-Appendix A Related Work
-A.1 VLM/LLM-based Agentic Frameworks in Embodied Tasks
-The rapid advancement of VLMs/LLMs has led to diverse agent frameworks in embodied environments [42, 33, 24].
-Embodied tasks involve partial observability and long-horizon planning, requiring memory systems to retain context.
-Some use time-ordered context buffers for short-term memory (due to VLMs/LLMs’ limited long-context processing) [42, 29]; others adopt experience buffers as long-term semantic memory [14, 32].
-For long-duration tasks, skill libraries serve as procedural memory, with agents accumulating skills via interaction [38, 35].
-However, in real-world settings, the low-level executor may fail to complete the task, making it challenging to construct a reusable, code-based skill library.
-So, explicit procedural memory still needs to be improved in real-world settings.
-Moreover, Recent efforts integrate diverse memories [44, 35, 1] but focus on virtual/GUI environments, leaving real-world multi-modal memory support for long-term planning under-explored.
-A.2 Vision Language Action Model
-Current work on VLA models uses imitation learning to output low-level controls from language and visuals [5, 47, 4, 22] but is limited to tabletop tasks and single actions, restricting long-horizon planning.
-VLAs lack long-term execution abilities, while high-level agents excel at planning.
-Recent works combine high-level frameworks with VLA executors, some augmented with simple memory [31, 34, 43, 41] for longer tasks.
-However, real-world robots need more sophisticated memory to handle continuous multi-task operations over extended periods.
-A.3 Memory Frameworks
-Many previous works improve long-term planning via memory systems: Voyager [38] uses a skill library in Minecraft but lacks diverse memory types; CoELA [44] includes procedural, semantic, and episodic memory with a task-specific 2D map; MSI-Agent [14] utilizes insight as long-term memory for in-task learning.
-Hippo Retrieval Augmented Generation (RAG) [18] mimics the hippocampus and introduces KGs as long-term memory indices [6, 9], enhancing retrieval.
-However, the previous approach is mainly focused on constructing a KG with a static long context, such as a book, but it is hard to update the graph.
-We need to update the information in KG for the embodied task.
-Our approach builds a more general LLM-based memory system using a dynamic KG like Hippo RAG, which is designed for embodied tasks.
-Furthermore, we summarize the differences among different memory systems in previous work.
-The comparison is shown in the Table 3.
-Table 3: Comparison of Memory-Related Methods in Embodied Agent
-Method Multimodal Episodic Semantic Spatial Temporal Procedural Memory Implementation Real Robot
-NeSyC [12] Symbolic logic rules
-Reflexion [32] Buffer
-Voyager [38] RAG
-MSI-Agent [14] Database, RAG
-CoELA [44] Top-down semantic map
-Cradle [35] RAG
-Agent-S [1] RAG
-Expel [46] Buffer
-AutoManual [8] Buffer
-HiRobot [31] /
-Being-0 [43] Buffer
-RoboOS [34] Scene graph, database
-RoboMemory (Ours) RAG,KG
-Appendix B Additional Experiments
-B.1 Error Analysis
-We summarize the common errors of RoboMemory in the previous experiments. We classify errors into three main types: planning errors, reasoning errors, and perception errors.
-The planning errors occur when the planner fails to generate correct actions.
-The reasoning errors occur when the planner and critic cannot properly process input information (including current observations and memory), even when the input is correct.
-Perception errors occur when incorrect information is provided to the planner-critic module.
-Figure 6: The reason why RoboMemory failed to complete the task.
-We analyze RoboMemory trajectories for failed tasks.
-We identify error types based on the above definitions.
-A single task may contain multiple errors.
-We calculate the occurrence probability of each error type to show RoboMemory’s strengths and weaknesses.
-The results are shown in Figure 6.
-We can observe that among all error types, the planning errors are the most common.
-This means that even though the memory modules can provide comprehensive information about the RoboMemory agent’s previous experience and spatial and temporal memory for the current task, the planner module may still not provide good action plans.
-This may be due to the capability of the pretrained base model.
-The most common perception error is the hallucination error.
-We can observe that although some hallucinations can be handled by the critic module or memory information, there are still some cases in which the planner ignores all insights from memory and critic and fails to complete the task.
-The detailed examples and discussions are provided in Appendix E.
-B.2 Evaluate on EB-Habitat
-Figure 7: Comparison of Success Rates (SR) and Goal Condition Success Rates (GC) across difficulty levels between RoboMemory and baseline methods on EB-Habitat.
-Similar to EB-ALFRED, we also deploy RoboMemory to EB-Habitat.
-Both EB-Habitat and EB-HALFRED are subsets of EmbodiedBench.
-We evaluate our model on the Base and Long subsets.
-Each subset contains 50 different trails.
-The results are shown in Figure 7.
-The results demonstrate that our RoboMemory can adapt well to different environments.
-It achieves significant improvements over the baseline across various settings.
-On average, the success rate increases by 24% compared to the SOTA Multi-Agent Method.
-The goal-conditioned success rate improves by 12%.
-These improvements indicate that RoboMemory enhances the agent’s embodied intelligence in various environments. The key factor is its complete memory system.
-B.3 Additional Efficiency analysis
-We analyze the evolution of the spatial KG during long trajectories in EB-ALFRED, focusing on the first 20 iterations (with 95% confidence intervals).
-As shown in Figure 8, the total number of spatial relationships in the KG (red line) increases gradually over iterations as RoboMemory is exploring the environment.
-In contrast, the number of relationships retrieved for update at each iteration (blue line) remains relatively stable, typically ranging around 10 edges per iteration.
-This stability is achieved because our method only updates a local subgraph relevant to the current observation.
-We define the retrieval ratio as the proportion of relationships updated at each iteration relative to the total number of relationships in the KG.
-As shown in Figure 8, this ratio (illustrated by gray bars) decreases steadily from 76% initially to 28% at iteration 20.
-This trend indicates that, as the KG grows, each update affects a progressively smaller fraction of the entire graph.
-This demonstrates that our spatial KG update mechanism effectively localizes modifications, ensuring computational efficiency and mitigating interference through context-aware incremental updates.
-Figure 8: Average relationships related to update in spatial memory in each step.
-Appendix C Dynamic Spatial Memory Update Algorithm
-C.1 Detailed Algorithm of Spatial KG Update
-Algorithm 1 Retrieval-based Incremental Knowledge Graph Update Algorithm
-0: New spatial knowledge graph $G_{\text{new}}=(V_{\text{new}},E_{\text{new}})$, main spatial knowledge graph $G=(V,E)$, queries $q\in Q$, entity & query embeddings $\mathcal{E}:V\cup Q\rightarrow\mathbb{R}^{d}$, maximum number of retrieved vertices $n$, maximum k hops $k$, vlm-base conflict resolver $\text{ResolveConflict}(\cdot)$
-0: Updated consistent knowledge graph $G^{\prime}$
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-9: Remove isolated vertices from $G^{\prime}$
-10:return $G^{\prime}$
-Spatial Memory is a dynamically updated KG-based module designed to overcome agents’ limitations in spatial reasoning.
-Specifically, our Spatial Memory is formulated as a directed KG $G=(V,E)$, where $V$ denotes the set of all objects in the environment.
-Each object is a vertex in the KG.
-Each object’s name is encoded into a single semantic embedding vector via a pretrained embedding model [45].
-The edge set $E$ captures spatial relationships between objects, each represented as a triple (e.g., $obj_{1},\text{relationship},obj_{2}]$).
-To update the Spatial KG, we need to continuously extract new relationships from current observations and update the old relationships in the Spatial KG.
-We use a VLM-based conflict resolver to address this problem.
-However, the more relationships provided to the conflict resolver, the more time it needs to update the KG.
-So we need to update relationships that are only related to the current situation.
-We design an algorithm that retrieves a sub-graph of KG that includes all vertices related to the current situation and both old and new relationships among them.
-We provide the sub-graph and new relationships to the conflict resolver.
-We need the conflict resolver to update the sub-graph based on information from the new relationships.
-The algorithm is shown in Algorithm 1.
-To update $G$, we make use of the information provided by the step summarizer and query generator from the information preprocessor introduced in Sec.
-2.1.
-First, a pretrained VLM-based Relation Retriever extracts the latest spatial relationships $G_{\text{new}}=(V_{\text{new}},E_{\text{new}})$ from the information provided by the step summarizer, which records high-level information in the current observation.
-Next, natural language queries (represented as $q\in Q$) (provided by the query generator) are used to retrieve relevant object vertices from $G$ via cosine similarity search.
-We select the top $n$ similar vertices compared with the query.
-These vertices are represent by $V_{\text{similar}}$.
-Spatial KG maintains the relationships among different objects, so if we want to retrieve spatial information from spatial KG, we need to search for other objects that are related to the objects we observed in the current observation.
-In this way, we not only remember objects we can see, but also know the spatial information of the objects we cannot see.
-So we choose the k-hop algorithm to expand $V_{\text{similar}}$ using a K-hop neighborhood algorithm to capture contextually related objects.
-The K-hop algorithm is represented as $\text{K-hop}_{k}(V,G)$, which returns all vertices reachable within $\leq k$ hops from any vertices in $V$.
-The retrieved vertices are $V_{\text{expand}}$.
-We combine the vertices retrieved by cosine similarity and their K-hop neighbors to $V_{\text{retrieved}}$.
-However, we need to resolve the conflict between new and old relationships.
-So $V_{\text{retrieved}}$ and the relationships among $V_{\text{retrieved}}$ is not enough.
-We need new vertices and relationships involved in the graph we provided to the VLM-based conflict resolver.
-To extract all vertices and relationships for the VLM-based conflict resolver and relationships, we not only need the relationships from KG (old information) and $G_{\text{new}}$, which represent new information.
-We need to connect old information and new information.
-To achieve this goal, we merge $G_{\text{new}}$ to $G$, which add new edges and vertices to $G$.
-We denote the merged KG as $G_{\text{union}}$.
-In $G_{\text{union}}$, we mix out-of-date and latest information.
-Then, we extract an induced sub-graph of $V_{\text{merged}}=V_{\text{retrieved}}\cup V_{\text{new}}$ from $G_{\text{union}}$.
-As both vertices from old graph $G$ and new graph is mixed in $V_{\text{merged}}$ and both edges from old and new graph is in $G_{\text{union}}$, the retrieved induced sub-graph $G_{\text{local}}$ contains all out-of-date and latest relationships among vertices that is related to current situation.
-As $G_{\text{local}}$ contains all out-of-date and latest relationships and those relationships may have conflicts, a VLM-based conflict resolver (represent as $\text{ResolveConflict}(\cdot)$) is designed to resolve conflicts in $G_{\text{local}}$, and make sure that the relationship is the latest.
-The conflict resolver will take in $G_{\text{new}}$ and $G_{\text{local}}$, where $G_{\text{local}}$ is the graph waiting for update and $G_{\text{new}}$ provide update signal.
-The VLM-based conflict resolver will perform necessary updates such as adding vertices or inserting, deleting, or modifying relationships in $G_{\text{local}}$ based on $G_{\text{new}}$.
-The reconciled subgraph is then merged back into $G$, and any vertices that have lost all connections to other vertices during the update are pruned.
-This design offers two key advantages: (1) Efficiency via localized updates: By restricting modifications to a context-relevant subgraph, we significantly reduce the number of relationships processed per update.
-Since VLMs struggle with reasoning over large sets of relationships, this constraint substantially improves both the efficiency and effectiveness of VLM-based KG updates.
-(2) Dynamic adaptability: The system continuously maintains up-to-date spatial knowledge, enabling agents to operate robustly in dynamic real-world environments.
-Figure 9: Visualization of Spatial Memory’s dynamic update process.
-C.2 Example of Dynamic Spatial Memory Update process
-In RoboMemory’s Spatial Memory, the KG is dynamically constructed during environment exploration.
-As illustrated in Figure 9, we demonstrate the progressive expansion of the KG in Spatial Memory as the agent navigates through the environment.
-The figure indicates a continuous growth in the number of both vertices and edges of the KG as exploration progresses.
-Notably, the KG undergoes dynamic updates through RoboMemory’s environmental interactions.
-For example, the initial KG state displays the relation “I am near the apple.
-But as the agent picks up the apple in the third step, in the fourth KG, the relationship becomes “I hold the apple”.
-This demonstrates RoboMemory’s capability for dynamic KG maintenance and expansion.
-By querying this KG, the Planner-Critic module gains access to rich spatial information, empowering RoboMemory with robust spatial memory capabilities that significantly enhance its performance in both TextWorld and EmbodiedBench environments.
-C.3 Proof of Dynamic Spatial Memory Update Algorithm
-Theorem 1 (Upper Bound on K-hop Vertex Extraction in Directed Graphs).
-Let $G=(V,E)$ be a finite directed graph with maximum out-degree $D\geq 1$, and let $\mathcal{S}\subseteq V$ be a set of $M$ source vertices.
-Define the K-hop neighborhood $\mathcal{N}_{K}(s)$ of a vertex $s\in\mathcal{S}$ as the set of vertices reachable from $s$ via directed paths of length at most $K$.
-Then the total number of distinct vertices in the union of all K-hop neighborhoods,
-$\mathcal{N}_{K}(\mathcal{S})=\bigcup_{s\in\mathcal{S}}\mathcal{N}_{K}(s),$
-Satisfies the following upper bound:
-$|\mathcal{N}_{K}(\mathcal{S})|\leq\begin{cases}M\cdot\dfrac{D^{K+1}-1}{D-1},&\text{if }D>1,\\ M\cdot(K+1),&\text{if }D=1.\end{cases}$
-Proof.
-For any vertex $s\in\mathcal{S}$, the number of distinct vertices reachable from $s$ within $i$ hops is at most $D^{i}$, assuming the worst-case scenario where each vertex encountered has the maximum out-degree $D$, and all neighbors are distinct and non-overlapping.
-Thus, the size of the K-hop neighborhood of a single vertex satisfies:
-$|\mathcal{N}_{K}(s)|\leq\sum_{i=0}^{K}D^{i}=\begin{cases}\dfrac{D^{K+1}-1}{D-1},&\text{if }D>1,\\ K+1,&\text{if }D=1.\end{cases}$
-Since there are $M$ such source vertices and assuming no overlaps between their K-hop neighborhoods (worst case), the union size satisfies:
-$|\mathcal{N}_{K}(\mathcal{S})|\leq M\cdot|\mathcal{N}_{K}(s)|.$
-Substituting the bound on $|\mathcal{N}_{K}(s)|$ gives the result. ∎
-Theorem 2 (Upper Bound for K-hop Vertex Extraction in Normalized Directed Graphs).
-Let $G=(V,E)$ be a finite directed graph with $|V|=n$ vertices.
-Assume the maximum out-degree is at most $D_{\max}=Dn$, and the maximum in-degree is at most $N_{\max}=Nn$, where $D,N\in(0,1]$ are constants.
-Let $\mathcal{S}\subseteq V$ be a set of $M$ source vertices.
-Define $\mathcal{N}_{K}(\mathcal{S})$ as the union of all vertices reachable from $\mathcal{S}$ via paths of length at most $K$, using only outgoing edges.
-Then the number of extracted vertices satisfies:
-$|\mathcal{N}_{K}(\mathcal{S})|\leq\min\left\{n,\;M\cdot\frac{(Dn)^{K+1}-1}{Dn-1}\right\}.$
-In particular, when $Dn\gg 1$, we have the approximation:
-$|\mathcal{N}_{K}(\mathcal{S})|\lessapprox M\cdot(Dn)^{K}.$
-Proof.
-For each vertex $s\in\mathcal{S}$, the maximum number of reachable vertices within $i$-hops is at most $(Dn)^{i}$ under the assumption of maximum out-degree and no overlap.
-Summing over hops from 0 to $K$, we get for each root:
-$|\mathcal{N}_{K}(s)|\leq\sum_{i=0}^{K}(Dn)^{i}=\frac{(Dn)^{K+1}-1}{Dn-1}.$
-Assuming no overlap among the $M$ source vertex expansions (worst case), we have:
-$|\mathcal{N}_{K}(\mathcal{S})|\leq M\cdot\frac{(Dn)^{K+1}-1}{Dn-1}.$
-Since the total number of vertices in the graph is $n$, this quantity is also trivially bounded above by $n$, yielding the result. ∎
-Appendix D Additional Environment settings
-D.1 EB-ALFRED and EB-Habitat
-We adopt the same environment parameters as in EmbodiedBench.
-The maximum steps per task are set to 30, with image inputs of size 500 $\times$ 500.
-The temporal memory buffer length is set to 3.
-However, we modified the action formats of EB-ALFRED and EB-Habitat to simulate real-world scenarios better.
-Specifically, we define different action APIs (Python functions), where each action takes an object parameter indicating its target.
-We extract all possible objects from the environment as inputs to the Agent.
-The Agent must select appropriate actions and object parameters based on task requirements.
-Compared to the original interaction method in EmbodiedBench (which enumerates all possible actions, including both action names and target objects, and requires the Agent to choose), our approach offers greater flexibility.
-The detailed action APIs are presented in Table 4.
-Since EB-ALFRED and EB-Habitat provide comprehensive high-level action APIs, we do not employ the VLA-Based Low-Level Executor in these environments.
-Instead, we utilize the built-in low-level controllers from EmbodiedBench.
-D.2 Real-world experiments
-Table 4: Robot Action Command For different environments
-Action Type EB-ALFRED EB-Habitat Real World
-Navigation find(obj) navigate(point) navigate_to(point)
-Pick Up Object pick_up(obj) pick(obj) pick_up(obj)
-Drop to Ground drop()
-Place to Receptacle put_down() place(rec) put_down_to(rec)
-Open Object open(obj) open(obj) open(obj)
-Close Object close(obj) close(obj) close(obj)
-Turn On turn_on(obj) turn_on(obj)
-Turn Off turn_off(obj) turn_off(obj)
-Slice Object slice(obj)
-Task Complete task_complete()
-We construct a common kitchen scenario to evaluate the RoboMemory framework’s interactive environmental learning capabilities in real-world settings.
-Using Mobile ALOHA [15] as our physical robotic platform, we design three categories of tasks: (1) Pick up & put down: The agent must locate a specified object among all possible positions and place it at a designated location.
-This task tests the model’s basic object-searching and planning abilities.
-(2) Pick up, operate & put down: Building upon the first task, the agent must additionally perform operations such as heating or cleaning the object.
-This task requires longer-term planning, which is crucial in embodied environments.
-(3) Pick up, gather & put down: The agent must place specified objects into a movable container and then move the container to a target location.
-This task evaluates the agent’s understanding of object relationships, requiring it to remember the positions of at least two objects (the container and the target item) and their spatial relationship.
-For each type of task, we design 5 tasks.
-So our experiments include 15 long-term real-world tasks.
-To adapt to the real-world setup, we define high-level action APIs similar to those in EB-ALFRED and EB-Habitat.
-Additionally, we train a VLA-based model to execute tasks according to our action APIs.
-The detailed action APIs are presented in Table 4.
-For the low-level executor, we use one main camera and two arm-mounted cameras as input, each with a resolution of 640 $\times$ 480. The temporal memory buffer length is set to 3.
-In our experiments, we set the maximum steps per task to 25.
-We also provide an API for actively terminating tasks.
-Since real-world environments lack direct success/failure feedback, RoboMemory must autonomously determine task completion.
-To prevent excessively long task execution, we enforce termination after 25 steps if no success is achieved.
-A single main camera (640 $\times$ 480 resolution) records video during action execution as input for RoboMemory’s higher-level processing.
-Table 5: Dataset statistics and training hyperparameters for robotic manipulation tasks.
-Dataset Statistics Training Configuration
-Action Type #Episodes Parameter Value
-Turning on/off faucet 142 Optimizer AdamW
-Picking up & Placing basket on counter 63 Batch size $32\times 6$
-Picking up & Placing basket in sink 72 Training steps 10,000
-Picking up & Placing banana into basket 114 Learning rate $6.12\times 10^{-5}$
-Throwing bottle into trash bin 132 warm up step 500
-Placing gum box on dish 120 LoRA Configuration
-Picking up & Placing cup on plate 51 rank 16
-Picking up & Placing dish into sink 69 $\alpha$ 16
-Throwing paper ball into trash bin 135 Resource Usage
-Open/close oven 142 GPU A100-80GB $\times$ 6
-Total episodes 1040 Training time 12 hours
-D.3 Training Details of Low-Level Executor
-We use the $\pi_{0}$ model as our foundation model.
-We collected 1,040 data samples over 10 types of tasks for fine-tuning.
-We use LoRA fine-tuning to save resources during fine-tuning.
-The specific fine-tuning parameters and action types are given in Table 5.
-For tasks involving both pick-up and place actions, we split these tasks into separate pick-up and place actions.
-These are then treated as two distinct data samples during training.
-The separation of pick-up and place action allows the VLA to carry an object in its hand.
-For training, we used a server with six A100-80GB GPUs.
-The total training time was 12 hours.
-Besides, we use the built-in LiDAR SLAM system of the Mobile ALOHA robot base as the navigation action actuator.
-We define five typical navigation points, similar to EB-Habitat.
-We used SLAM to navigate between these navigation points.
-D.4 Hyperparameters of RoboMemory
-In this section, we describe the hyperparameter settings for the upper brain of RoboMemory: the information preprocessor and the Comprehensive Embodied Memory.
-Importantly, we use a unified set of hyperparameters across all experimental settings, including EB-ALFRED, EB-Habitat, and real-world deployments.
-The information preprocessor consists of two components: a step summarizer and a query generator.
-Given multimodal inputs at each step, the step summarizer produces a single natural language description, while the query generator concurrently formulates 454-5 distinct natural language queries.
-The Comprehensive Embodied Memory integrates four memory modules: Temporal, Spatial, Semantic, and Episodic Memory.
-The Temporal Memory is implemented as a fixed-size buffer with a maximum capacity of $4$ entries.
-For Spatial Memory, during similarity-based retrieval, we first identify the top $N=3$ most relevant vertices and then perform a K-hop graph traversal with $K=2$.
-The Episodic Memory retrieves the top $N=5$ most relevant past experiences for each query.
-The Semantic Memory maintains hierarchical summaries at both the action and task levels; during retrieval, it returns $N_{s}=2$ action-level and $N_{t}=2$ task-level summaries.
-Furthermore, memory updates (e.g., insertion, modification, or deletion) are applied only to the top $N_{\text{update}}=10$ most relevant entries in the Semantic Memory to ensure efficiency and coherence.
-Appendix E Supplementary Examples for Qualitative Analysis
-E.1 Real World
-Figure 10: Case that a task is failed, but the experience can help RoboMemory to succeed in the next try.
-In Figure 10, we demonstrate an example of RoboMemory learning through trial and error in a real-world environment.
-Our task is “place a banana into the oven.” This task required RoboMemory to complete the objectives of finding the banana, picking it up, and transporting it to the oven.
-We observed that RoboMemory became stuck in an infinite loop during the first attempt.
-The banana was randomly placed on the “kitchen counter,” but RoboMemory overlooked this navigation target and remained trapped, exploring other navigation targets instead.
-However, based on this bad attempt, the semantic memory summarized that the robot should not repeatedly search in locations where the “banana” could not be found.
-Meanwhile, the episodic memory recorded what RoboMemory had done and the outcomes during the first attempt.
-Based on the information provided by semantic and episodic memory, in the second attempt, RoboMemory recognized that it had not previously tried navigating to the “kitchen counter.” After attempting this, it successfully completed the task.
-This example illustrates the role of RoboMemory’s long-term memory.
-Figure 11: Case that a task is successful.
-We also provide an example that completes the task in the first attempt.
-The example is shown in Figure 11.
-This example demonstrates that the RoboMemory has the ability to handle some relatively complex tasks in the real world.
-The task in this example is “Place a box of gum into the basket and put the basket on the kitchen counter”.
-Because two objects in different positions are involved in this task, RoboMemory has to memorize the position of at least one object to achieve the goal.
-With the help of the spatial memory, RoboMemory completes the task successfully.
-E.2 EB-ALFRED
-We select three examples in EB-ALFRED to show the errors that RoboMemory may encounter and the reasons why or why not RoboMemory can achieve the goal.
-E.2.1 Successful example
-Figure 12: Case that a task is successful with the help of the critic and spatial memory modules.
-We select a successful example to show how RoboMemory performed in the EB-ALFRED environment. The example trajectory is shown in Figure 12.
-The task of this example is “set a plate with a spoon on it on the kitchen table”.
-However, in step 10, the Planner seems to ignore the temporal information from memory modules.
-RoboMemory thinks that it still needs to pick up the spoon (even though it has already placed a spoon in the plate).
-However, with the help of the critic, it finally becomes aware that picking up another spoon is redundant, so RoboMemory goes back to the current trajectory and successfully completes the task at the end.
-In this example, RoboMemory successfully overcame the hallucination and eventually achieved the goal. This example demonstrates that the critic module can help RoboMemory to overcome error cases.
-E.2.2 Failed example
-Figure 13: Case that a task fails in an infinite loop because the critic module failed to stop the agent when its planned action is no longer suitable.
-We demonstrate a representative example of the Critical Error.
-The example trajectory is shown in Figure 13.
-In this example, the task involves slicing and heating a tomato and moving the heated tomato slice to the trash can.
-Initially, RoboMemory successfully sliced the tomato with a knife.
-But when the planner plans the whole sequence, it forgets to drop the knife before picking up the tomato (this is necessary because in EB-ALFRED, the robot can only hold one object at a time).
-The critic and the planner should notice this situation and ask the critic to replan, as RoboMemory failed to pick up a tomato slice.
-However, the critic module ignores this issue, and thus, after it heats the knife instead of a tomato slice, it stacks in an infinite loop.
-Figure 14: Case that a task fails in an infinite loop because of inaccurate action planning.
-Besides, we provide another example demonstrating a representative failure caused by inaccurate action planning.
-The example trajectory is shown in Figure 14.
-In the trajectory, RoboMemory is asked to place two CDs into the drawer.
-However, at step 6, the robot failed to select correct CD object.
-In this experiment, RoboMemory has already put CD_2 into the drawer, but it keeps picking up CD_2 even though the memory has clearly indicated that CD_2 has already been put down.
-So we classify this as inaccurate action error.
-This indicates that the planner failed to comprehensively integrate information from both the memory and information-gathering modules, resulting in inaccurate action planning.